### v1

# Holographic entanglement negativity and thermodynamics in backreacted AdS black hole

Abstract
We investigate holographic entanglement negativity (HEN) as a probe of mixed state quantum correlations in a deformed AdS black hole background with backreaction sourced by a string cloud. The bulk geometry is dual to a strongly coupled large-$N_{c}$ gauge theory at finite temperature, backreacted by a finite density of heavy static fundamental quarks. Alongside holographic entanglement entropy (HEE), we analyze entanglement thermodynamics and establish a first law like relation in the small backreaction regime by identifying the deformed black hole at zero temperature as the natural ground state of the dual field theory. Using analytic expansions in the low and high temperature limits, we compute the HEN for adjacent, disjoint and bipartite subsystem configurations and examine its dependence on the backreaction parameter $\rho$. Our results reveal a clear regime dependence: at low temperature, backreaction suppresses the negativity, signaling a reduction of distillable quantum correlations while at high temperature the negativity is enhanced due to the monotonic $\rho$-dependence of near horizon contributions. We further compare the behavior of HEN with HEE and mutual information, demonstrating that entanglement negativity provides a sharper diagnostic of the competition between thermal and quantum correlations in holographic plasmas with matter backreaction.
1 Introduction
Strongly coupled quantum systems appear in a wide range of physical settings ranging from condensed matter systems and cold atomic gases to the quark–gluon plasma formed in heavy-ion collisions. A central difficulty in analyzing such systems is the characterization of their quantum correlations, particularly in regimes where perturbative methods fail. The AdS/CFT correspondence [Maldacena:1997re, Witten:1998qj, Gubser:1998bc] offers a powerful non-perturbative framework by relating strongly coupled gauge theories to weakly coupled gravitational theories in higher dimensional spacetimes. Within this holographic framework, quantum information theoretic measures such as entanglement entropy, mutual information, and computational complexity have emerged as valuable probes of the correlation structure of the dual field theories. Among these, the holographic computation of entanglement entropy through the Ryu–Takayanagi (RT) prescription and its covariant generalization [Ryu:2006bv, Hubeny:2007xt] has played a central role. The RT prescription has found extensive application in investigating the entanglement entropy of conformal field theories at finite temperature which are holographically dual to AdS-Schwarzschild black holes [Cadoni:2010ztg, Solodukhin:2006xv, Hubeny:2012ry, Fischler:2012ca, Fischler:2012uv]. Detailed analyses of the resulting low and high temperature behaviors of holographic entanglement entropy led to the formulation of a thermodynamics like first law for the entanglement entropy in [Bhattacharya:2012mi, Allahbakhshi:2013rda]. Entanglement thermodynamics has also been investigated for charged black holes in AdS4 backgrounds. In particular, the effects of temperature and charge on the entanglement entropy of the boundary field theory dual to charged AdS black holes were studied in [Chaturvedi:2016kbk], with extensions to higher-dimensional charged black holes presented in [Karar:2018ecr]. Furthermore, entanglement thermodynamics has been explored in a variety of other scenarios including non-conformal and non-relativistic geometries [Blanco:2013joa, Dey:2014voa, Mansoori:2015sit, Park:2015afa, Chakraborty:2014lfa, Mishra:2015cpa, McCarthy:2017amh, OBannon:2016exv, Sun:2016til, Nadi:2019bqu, Saha:2019ado, Maulik:2020tzm].
The entanglement entropy does not provide a faithful characterization of quantum correlations in mixed states as it captures both quantum and classical contributions. To extract the purely quantum correlations in such states, entanglement negativity was introduced in quantum information theory [Vidal:2002zz, Plenio:2005cwa]. Entanglement negativity serves as computable upper bound on distillable entanglement, and monotonic under local operations and classical communication (LOCC). Further it vanishes for separable states, thereby making it a robust measure of quantum correlations in mixed state scenarios. In recent years, entanglement negativity has gained prominence as a powerful diagnostic of quantum correlations in quantum field theories, capturing information inaccessible through entanglement entropy alone. A field-theoretic formulation based on the replica trick was developed in [Calabrese:2012nk] enabling the computation of negativity in extended systems. This approach was later applied to finite temperature conformal field theories [Calabrese:2014yza] and to lattice models such as the critical Ising chain [Calabrese:2013mi]. The time-dependent behavior of entanglement negativity have also been explored in both local and global quantum quenches revealing universal features of entanglement spreading in conformal field theories [Wen:2015qwa, Coser:2014gsa]. These results collectively establish entanglement negativity as a central tool in characterizing quantum correlations in many body systems. Motivated by these developments, several holographic prescriptions for entanglement negativity have been proposed [Rangamani:2014ywa, Chaturvedi:2016rcn, Chaturvedi:2017znc, Jain:2017aqk, Jain:2017uhe, Malvimat:2018txq, Malvimat:2018ood, Chaturvedi:2016rft, Jain:2017xsu, Jain:2018bai, KumarBasak:2020viv, Mondal:2021kzj, Afrasiar:2021hld, Kudler-Flam:2018qjo, Kusuki:2019zsp, KumarBasak:2020eia, KumarBasak:2021lwm] extending the intuition gained from the RT prescription. The HEN turns out to be an effective measure to study the behavior of entanglement in the vicinity of critical points for strongly coupled gauge theory [Karan:2023hfk]. These studies reveal that negativity is sensitive to critical behavior and captures features that are not visible through entanglement entropy alone. In parallel, HEN provides a robust characterization of quantum correlations in the presence of thermal and deformation effects in quantum field theories at finite temperature [Pant:2024eno]. Together, these developments underscore the effectiveness of HEN as a better probe of mixed state entanglement in holographic frameworks. Closely related insights arise from the study of reflected entropy, that offers an alternative mixed state measure based on the entanglement structure of canonical purifications [Dutta:2019gen]. While reflected entropy captures both quantum and classical correlations, its holographic dual via the entanglement wedge cross section provides complementary geometric information that enriches the understanding of correlation structure beyond what is accessible through negativity alone. Together, holographic entanglement negativity and reflected entropy describes a broad framework for probing mixed state entanglement in strongly coupled dual field theories.
Recent studies have highlighted the significance of investigating entanglement measures in bulk geometries that include matter backreaction as such effects capture additional physical features of the dual strongly coupled field theories. An instructive example is provided by a deformed AdS black hole geometry sourced by a homogeneous string cloud [Chakrabortty:2011sp] which is holographically dual to a large-$N_{c}$ gauge theory at finite temperature backreacted by a uniform density of static heavy quarks. The backreaction parameter encodes the quark density and modifies both the near boundary and near horizon geometry, providing a controlled setup to investigate how heavy-flavor matter influences entanglement structure, thermodynamics, and mixed-state quantum correlations. This setup has been extensively employed to explore transport properties, entanglement structure and chaotic behavior in the dual field theory [Chakrabortty:2016xcb, Chakrabortty:2020ptb, Chakrabortty:2022kvq, Jain:2023xta].
A natural question that arises is whether holographic entanglement entropy admits a relation similar to the first law of thermodynamics in such backreacted field theories. To address this, we study entanglement thermodynamics in deformed strongly coupled field theories in the context of AdS4/CFT3. Specifically, we analyze the dependence of the entanglement entropy on temperature and backreaction for a strip shaped subsystem in the boundary field theory dual to a deformed AdS4 black hole. Furthermore, while the impact of backreaction on entanglement entropy and related probes has been widely investigated, the behavior of HEN in this background remains largely unexplored. Given that entanglement negativity provides a better probe of distillable quantum entanglement in mixed states, a systematic analysis in this context is both natural and timely. In this work, we present a detailed analysis of HEN for adjacent, bipartite and disjoint subsytems in the the string cloud deformed AdS black hole geometry. Employing analytic expansions in both low and high temperature regimes, we find a clear dichotomy in the effect of backreaction. At low temperature, the leading $\rho$-dependent correction enters as $(1-\rho)T_{f}^{\,d}$ and suppresses the HEN which indicates the inclusion of heavy static quarks enhances thermal correlations at the cost of quantum correlations. By contrast, in the high temperature regime, the HEN scales with $T_{f}$ and is enhanced with increasing $\rho$, as the near horizon contributions encoded in the $\widetilde{S}(\rho,d)$ coefficients grow monotonically with the quark density. This dichotomy persists for adjacent and disjoint intervals while in the bipartite configuration, the interplay between $\rho T_{f}^{\,d-1}$ and $(1-\rho)T_{f}^{d}$ terms leads to richer, parameter-dependent behaviour. A comparison with HEE shows that while, HEE is dominated by thermal scaling and is insensitive to the separation of quantum and classical correlations, HEN captures the genuine quantum contribution and thus provides a sharper diagnostic of backreaction effects in strongly coupled plasmas. Further, HEN for two disjoint subsystems allow us to compute the critical separation at which the distillable entanglement vanishes. We give a master equation that can be used to find the critical separation for particular values of backreaction and dimensions.
The structure of the paper is as follows. In section˜2, we introduce the string cloud deformed AdS black hole geometry and its holographic dual interpretation. Section 3 provides a review of holographic entanglement entropy in this background, with particular emphasis on the analytic behavior in the low and high temperature regimes. In section˜4, we analyze the holographic entanglement entropy in boundary field theories dual to deformed AdS4 black holes at both zero and finite temperatures in the regime of small backreaction, and and establish a first law of entanglement thermodynamics at low temperatures. In section˜5, we present the computation of holographic entanglement negativity for adjacent subsystems followed by the analysis of bipartite configuration in section˜6. The HEN for the configuration of two disjoint subsystems is examined in section˜7 followed by a critical separation analysis in section˜8. Finally, in section˜9, we summarize our findings.
2 Background
We begin by briefly reviewing the holographic setup corresponding to a strongly coupled gauge theory with a finite density of heavy static fundamental quarks, which is dual to a deformed AdS black hole geometry. This deformation arises due to the backreaction of a homogeneous distribution of long, non-interacting, and static strings stretched along the radial direction in the bulk spacetime. The endpoints of these strings are attached to the AdS boundary and represent static fundamental quarks, while the string bodies encode the associated gluonic degrees of freedom [Chakrabortty:2011sp].
The gravitational sector of the dual theory is described by the following $(d+1)$-dimensional Einstein-Hilbert action with a negative cosmological constant:
$\mathcal{S}=\frac{1}{4\pi G_{d+1}}\int dx^{d+1}\sqrt{g}\left(\mathcal{R}-2\Lambda\right)+\mathcal{S}_{M},$ (2.1)
where the matter action $\mathcal{S}_{M}$ captures the effect of the string distribution and is given by
$\mathcal{S}_{M}=-\frac{1}{2}\sum_{i}\mathcal{T}_{i}\int d^{2}\xi\sqrt{-h}h^{\alpha\beta}\partial_{\alpha}X^{\mu}\partial_{\beta}X^{\nu}g_{\mu\nu}.$ (2.2)
Here, $g_{\mu\nu}$ denotes the $(d+1)$-dimensional spacetime metric, $h_{\alpha\beta}$ is the intrinsic worldsheet metric, and $\mathcal{T}_{i}$ represents the tension of the $i^{\text{th}}$ string. The corresponding Einstein’s equations, get modified by the energy-momentum tensor sourced by the strings:
$\mathcal{R}_{\mu\nu}-\frac{1}{2}\mathcal{R}g_{\mu\nu}+\Lambda g_{\mu\nu}=8\pi G_{N}T_{\mu\nu},$ (2.3)
where the stress-energy tensor $T^{\mu\nu}$ associated with the string cloud is
$T^{\mu\nu}=-\sum_{i}\mathcal{T}_{i}\int{d^{2}\xi\frac{1}{\sqrt{\mid{g_{\mu\nu}}}}\sqrt{-h^{\alpha\beta}}h^{\alpha\beta}\partial_{\alpha}X^{\mu}\partial_{\beta}X^{\nu}g_{\mu\nu}\delta_{i}^{d-1}(x-X_{i})}.$ (2.4)
Considering all the strings are identical with tension $\mathcal{T}$ and adopt the static gauge choice $t=\xi^{0}$, $r=\xi^{1}$ gives the following number density of the strings
$b(x)=\mathcal{T}\sum_{i=1}^{N}\delta_{i}^{(d-1)}(x-X_{i}),\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{with}\penalty 10000\ b>0,$
where $N$ denotes the total number of strings. Assuming a uniform distribution over a spatial volume $V_{d-1}$, the average density becomes:
$\tilde{b}=\frac{1}{V_{d-1}}\int b(x)d^{d-1}x=\frac{\mathcal{T}N}{V_{d-1}}.$
As suggested in [Chakrabortty:2011sp], the thermodynamic limit ($V_{d-1}\to\infty$), to keeps $N/V_{d-1}$ finite for maintaining constant quark density as $N\to\infty$. The only non-vanishing components of the energy-momentum tensor are,
$T_{00}=-\frac{\tilde{b}}{r^{3}}g_{tt},\qquad T_{rr}=-\frac{\tilde{b}}{r^{3}}g_{rr}.$ (2.5)
The suitable ansatz for solving modified Einstein equation in AdS is
$ds^{2}=-V(r)dt^{2}+\frac{dr^{2}}{V(r)}+\frac{r^{2}}{R^{2}}\delta_{ij}dx^{i}dx^{j},$ (2.6)
where the blackening factor $V(r)$ is defined as:
$V(r)=K+\frac{r^{2}}{R^{2}}-\frac{2m}{r^{d-2}}-\frac{2bR^{d-3}}{(d-1)r^{d-3}}.$ (2.7)
Here $m$ is the mass parameter and $K$ denotes the curvature of the spatial boundary and takes values 1-1, 0, or $1$ for hyperbolic, flat, or spherical geometries, respectively. In our analysis, we consider the planar ($K=0$) case only. Additionally, we define a dimensionless deformation parameter $b=\tilde{b}R^{d-1}$ for notational convenience. Transforming to the inverse radial coordinate $z=R^{2}/r$, the metric takes the following familiar form:
$ds^{2}=\frac{R^{2}}{z^{2}}\left(-h(z)dt^{2}+d\vec{x}^{2}+\frac{dz^{2}}{h(z)}\right),$ (2.8)
with the modified blackening function now expressed as
$h(z)=1-\frac{2m}{R^{2d-2}}z^{d}-\frac{2b}{(d-1)R^{d-1}}z^{d-1}.$ (2.9)
It is convenient to rewrite $h(z)$ in terms of the horizon location $z_{H}$ and a dimensionless quark density parameter $\rho$:
$h(z)=\left[1-\rho\left(\frac{z}{z_{H}}\right)^{d-1}+(\rho-1)\left(\frac{z}{z_{H}}\right)^{d}\right],$ (2.10)
where
$\rho=\frac{2bz_{H}^{d-1}}{(d-1)R^{d-1}}.$ (2.11)
Both $b$ and $\rho$ encode the information of the uniform string (quark) density in the bulk and on the boundary, respectively. These parameters will be used throughout the paper depending on the convenience. The Hawking temperature of this deformed AdS black hole is
$T=-\frac{1}{4\pi}\frac{d}{dz}h(z)\Biggr|_{z=z_{H}}=\frac{(d-\rho)}{4\pi z_{H}}.$ (2.12)
Positivity of the temperature imposes the constraint $0\leq\rho\leq d$. The lower limit corresponds to the undeformed case ($b=0$), while the upper bound corresponds to the extremal limit with vanishing Hawking temperature. The corresponding ADM mass of the black hole is related to the parameter $m$ via $M=\frac{(d-1)mV_{d-1}}{8\pi G_{d+1}}$, and the associated entropy density is:
$s=\frac{R^{d-1}}{4G_{N}^{d+1}}\frac{1}{z_{H}^{d-1}}.$ (2.13)
It is useful to introduce an effective temperature $T_{f}$, defined as:
$T_{f}=\frac{d}{4\pi z_{H}},$ (2.14)
which allows the entropy density to scale as $s\sim T_{f}^{d-1}$, a feature that becomes important in subsequent analyses of low temperature entanglement entropy.
The specific heat of the black hole computed from the variation of mass with temperature is given by
$C=\frac{\partial M}{\partial T}=\frac{V_{d-1}(d-1)R^{2d-2}(d(d-1)R^{2d}z_{H}-2bR^{4}z_{H}^{d})}{4G_{d+1}z_{H}^{d-1}(d(d-1)R^{2d}z_{H}+2(d-2)bR^{4}z_{H}^{d})}$ (2.15)
To ensure thermodynamic stability, the black hole must satisfy the condition:
$z^{{\rm min}}_{H}=R\left(\frac{d(d-1)}{2b}\right)^{\frac{1}{d-1}}.$ (2.16)
At zero temperature, the black hole acquires a negative minimum mass:
$m^{{\rm min}}=-\frac{R^{d-1}}{d}\frac{b}{z^{{\rm min}}_{H}}.$ (2.17)
Above this bound, the specific heat remains positive and continuous, indicating that the black hole configuration is thermodynamically stable. For $b>0$, black holes with a negative mass parameter $m$ can exist within the range $m^{\rm min}\leq m<0$. In contrast, when $b\leq 0$, black holes with negative $m$ do not possess real horizons. The corresponding free energy is given by:
${\cal F}=-\frac{(d-1)z_{H}+2bR^{2}(d-2)z_{H}^{d}}{16\pi z_{H}^{d+1}(d-1)}.$ (2.18)
To examine the dynamical (geometric) stability of the background spacetime, one can carry out a gauge-invariant analysis of gravitational perturbations using the formalism developed in [Ishibashi:2011ws, Kodama:2003jz, Ishibashi:2003ap, Kodama:2003kk, Kodama:2007ph]. Within this approach, the background spacetime is regarded as a product manifold:
$M^{2+p}={\mathcal{N}}^{2}\times{\mathcal{K}}^{p},$ (2.19)
Here, ${\mathcal{N}}^{2}$ denotes a two-dimensional spacetime with coordinates $(t,z)$, while ${\mathcal{K}}^{p}$ represents a $p$-dimensional maximally symmetric space spanned by the spatial coordinates $x^{i}$. Gravitational perturbations are classified into scalar, vector, and tensor modes on ${\mathcal{K}}^{p}$, which allows the Einstein equations to be analyzed in a decoupled manner. The tensor and vector sectors have been studied explicitly in [Chakrabortty:2011sp], where the background geometry was shown to be stable against these classes of perturbations.
3 Revisiting the holographic entanglement entropy and EWCS
Since entanglement thermodynamics and holographic entanglement negativity are primarily studied in this work, it is natural to first review the computation of holographic entanglement entropy in the backreacted AdS black hole geometry, as introduced in section˜2. Moreover, as reflected entropy is directly related to the entanglement wedge cross section, we also briefly review the EWCS. Both quantities were originally computed in [Chakrabortty:2020ptb].
3.1 Holographic entanglement entropy (HEE)
We now revisit the computation of HEE for a single strip-like region in the backreacted AdSd+1 black hole. This subsection provides all expressions later required for the computation of HEN. According to the RT prescription, HEE is given by
$S_{A}=\frac{\mathrm{Area}[\gamma_{A}]}{4G_{N}}\,$, (3.1)
where $\gamma_{A}$ is the extremal surface in the bulk anchored on the boundary of $A$. Consider a rectangular strip like subsystem $A$ with width $l$ along $x$ and infinite extent $L$ in other spatial directions defined by
$x_{1}\in\left[-\frac{l}{2},\,\frac{l}{2}\right],\quad x_{i}\in\left[-\frac{L}{2},\,\frac{L}{2}\right],\quad i=2,\dots,d-2\,.$ (3.2)
with $L\to\infty$. On a constant-time slice, the induced metric is
$ds^{2}=\frac{R^{2}}{z^{2}}\left[\frac{dz^{2}}{h(z)}+d{x_{1}}^{\,2}+d\vec{x_{i}}^{\,2}\right],$ (3.3)
where the blackening function $h(x)$ is given by eq.˜2.10. The profile of the extremal surface $z(x_{1})$ satisfies the RT area functional
$\mathcal{A}[z]=R^{d-1}L^{d-2}\int_{-l/2}^{l/2}dx_{1}\,\frac{1}{z^{d-1}}\sqrt{1+\frac{z^{\prime\,2}}{h(z)}}.$ (3.4)
Due to translational symmetry in $x_{1}$ direction, the conserved Lagrangian leads to
$\frac{1}{z^{d-1}}\frac{1}{\sqrt{1+\frac{z^{\prime 2}}{h(z)}}}=\frac{1}{z_{t}^{d-1}},$ (3.5)
where $z_{t}$ is the turning point of the minimal surface. Solving for $z^{\prime}(x_{1})$ and integrating gives the width and turning-point relation
$l/2=1/z_{t}^{d-1}\int_{0}^{z_{t}}z^{d-1}\left[h(z)\left(1-(z/z_{t})^{2d-2}\right)\right]^{-1/2}dz.$ (3.6)
Substituting (3.5) back into the area functional gives the area of the extremal surface and consequently the HEE is given by the RT formula as
$S_{A}=\frac{L^{d-2}R^{d-1}}{2G^{d+1}_{N}}\int_{0}^{z_{t}}\frac{dz}{z^{d-1}\sqrt{h(z)}\sqrt{1-\left(\frac{z}{z_{t}}\right)^{2d-2}}},$ (3.7)
To obtain the HEE in terms of boundary parameters, it is necessary to express the turning point $z_{t}$ as a function of the strip width and temperature. However, evaluating the integral in eq.˜3.6 for arbitrary dimension $d$ and finite temperature is analytically intractable. As discussed in [Chakrabortty:2020ptb], we therefore analyze the problem in two regimes i.e. the low and high temperature limits where the integrals can be computed explicitly, and subsequently use these results to extract the corresponding behavior of the HEE.
Low-temperature expansion:
By taking the low temperature limit i.e $T_{f}l\ll 1$ ( or $z_{t}/z_{H}\ll 1$) the expression of $z_{t}$ becomes
$z_{t}$ $=\frac{l}{2}\frac{\Gamma\left(\frac{1}{2d-2}\right)}{\sqrt{\pi}\,\Gamma\left(\frac{d}{2d-2}\right)}\Bigg[1-\frac{\rho}{2d(2d-2)}\left(\frac{\Gamma\left(\frac{1}{2d-2}\right)}{\Gamma\left(\frac{d}{2d-2}\right)}\right)^{2}\left(\frac{l}{2z_{H}}\frac{\sqrt{\pi}\,\Gamma\left(\frac{d}{2d-2}\right)}{\Gamma\left(\frac{1}{2d-2}\right)}\right)^{\!d-1}$
$\quad-\frac{1-\rho}{2(d+1)}\frac{\Gamma\left(\frac{2d}{2d-2}\right)\Gamma\left(\frac{1}{2d-2}\right)}{\Gamma\left(\frac{d}{2d-2}\right)\Gamma\left(\frac{1}{d-1}+\frac{1}{2}\right)}\left(\frac{l}{2z_{H}}\frac{\sqrt{\pi}\,\Gamma\left(\frac{d}{2d-2}\right)}{\Gamma\left(\frac{1}{2d-2}\right)}\right)^{\!d}+\cdots\Bigg].$ (3.8)
Now the corresponding area function is
$\mathcal{A}=\frac{2}{d-2}\frac{L^{d-2}R^{d-1}}{\epsilon^{d-2}}+R^{d-1}\left(\frac{L}{l}\right)^{\!d-2}S_{0}\left[1+\rho S_{2}\left(\frac{l}{z_{H}}\right)^{\!d-1}+(1-\rho)S_{1}\left(\frac{l}{z_{H}}\right)^{\!d}+\cdots\right].$ (3.9)
Here $S_{0},S_{1}$ and $S_{2}$ are terms that only depends on dimensions parameter $d$, and are given by
$\begin{split}S_{0}&=\frac{2^{d-2}\pi^{\frac{d-1}{2}}\Gamma{\left(\frac{-d+2}{2d-2}\right)}}{(d-1)\Gamma{\left(\frac{1}{2d-2}\right)}}\left(\frac{\Gamma{\left(\frac{d}{2d-2}\right)}}{\Gamma{\left(\frac{1}{2d-2}\right)}}\right)^{d-2}\\ S_{1}&=2^{-(d+1)}\pi^{-\frac{d}{2}}\frac{\Gamma{\left(\frac{1}{2d-2}\right)^{d+1}}}{\Gamma{\left(\frac{1}{d-1}+\frac{1}{2}\right)}\Gamma{\left(\frac{d}{2d-2}\right)^{d}}}\left[\frac{\Gamma{\left(\frac{1}{d-1}\right)}}{\Gamma{\left(\frac{-d+2}{2d-2}\right)}}+\frac{(d-2)}{(d+1)}\frac{2^{\frac{1}{d-1}}}{\sqrt{\pi}}\Gamma{\left(1+\frac{1}{2d-2}\right)}\right]\\ S_{2}&=2^{-d}\pi^{-(\frac{d-1}{2})}\left(\frac{\Gamma{\left(\frac{1}{2d-2}\right)}}{\Gamma{\left(\frac{d}{2d-2}\right)}}\right)^{d+1}\left[\frac{\Gamma{\left(\frac{d}{2d-2}\right)}}{\Gamma{\left(\frac{-d+2}{2d-2}\right)}}+\frac{(d-2)}{2d(d-1)}\right].\end{split}$ (3.10)
Finally the HEE in low temperature limit is given by
$S_{A}^{\mathrm{low}}=\frac{R^{d-1}L^{d-2}}{2G_{N}}\left[\frac{1}{(d-2)\epsilon^{d-2}}+S_{0}\,l^{2-d}+S_{0}S_{2}\,\rho\,l\left(\frac{4\pi T_{f}}{d}\right)^{d-1}+S_{0}S_{1}(1-\rho)\,l^{2}\left(\frac{4\pi T_{f}}{d}\right)^{d}+\cdots\right].$ (3.11)
The finite pieces $S_{0}\,l^{2-d}$, $S_{0}S_{2}\,\rho\,lT_{f}^{d-1}$, and $S_{0}S_{1}(1-\rho)l^{2}T_{f}^{d}$, are exactly the contributions that will appear in the HEN formulas.
In eq.˜3.11, the first term is divergent containing UV-cut off ($\epsilon$). In the low-temperature regime, the entanglement entropy exhibits nontrivial modifications due to backreaction. Unlike the charged plasma case [Kundu:2016dyk], where the leading correction scales as $T_{f}^{\,d}$, here the dominant term is proportional to $T_{f}^{\,d-1}$. This distinction arises from the blackening function $h(z)$: near the boundary the $z^{\,d-1}$ contribution from the quark density dominates over the mass term $\sim z^{d}$, making backreaction the leading source of corrections to the HEE at low effective temperature. It will be particularly interesting to examine whether a similar scaling behavior also manifests in the computation of HEN.
High-temperature expansion:
In the high temperature regime $T_{f}l\gg 1$ (or $z_{t}\to z_{H}$), the RT surface approaches the horizon. Although, the direct series expansions for $l$ and $A$ diverge in this limit, an appropriate combination of these quantities yields a finite and well behaved expression for the entanglement entropy as
$\mathcal{A}-\frac{L^{d-2}R^{d-1}}{z_{t}^{\,d-1}}\,l=\frac{2L^{d-2}R^{d-1}}{z_{t}^{\,d-2}}\int_{0}^{1}\frac{du}{\sqrt{h(u)}\,u^{\,d-1}\sqrt{1-u^{2d-2}}}-\cdots,$ (3.12)
where $u=\frac{z}{z_{t}}$. This leads to the finite part of HEE as
$S^{high}_{A}\simeq\frac{R^{d-1}}{4G_{d+1}}\left[V\left(\frac{4\pi T_{f}}{d}\right)^{\!d-1}\left(1+2\frac{d}{4\pi T_{f}l}\,\widetilde{S}(\rho,d)\right)\right],$ (3.13)
where $V=lL^{d-2}$ and
$\widetilde{S}(\rho,d)=\frac{\sqrt{\pi}\,\Gamma\left(-\frac{d-2}{2d-2}\right)}{(2d-2)\,\Gamma\left(\frac{1}{2d-2}\right)}+\int_{0}^{1}du\left[\frac{\sqrt{1-u^{2d-2}}}{u^{\,d-1}\sqrt{h(u)}}-\frac{1}{u^{\,d-1}\sqrt{1-u^{2d-2}}}\right].$ (3.14)
At high $T_{f}$, the leading term reproduces the thermal entropy density via its volume scaling, with $\rho$ entering only at subleading order.
Across both the low and high effective temperature limits, the quark density parameter $\rho$ plays a crucial role in determining the structure of the finite part of the entanglement entropy. At low $T_{f}$, the dominant correction scales as $T_{f}^{\,d-1}$, which originates from the backreaction-induced $z^{\,d-1}$ contribution through $h(z)$. In contrast, at high $T_{f}$ the leading term coincides with the thermal entropy density while the dependence on $\rho$ manifests only through subleading corrections.
3.2 Entanglement wedge cross section and reflected entropy
Unlike HEE, the Entanglement Wedge Cross Section (EWCS) required two non-overlapping intervals (say $A_{1}$ and $A_{2}$). It is a geometric quantity associated with the minimal bulk surface that divides the entanglement wedge of two boundary sub regions $A_{1}$ and $A_{2}$ into two parts homologous to each of them. The EWCS has emerged as an important mixed state correlation measure in holography, conjectured to be dual to the entanglement of purification [Takayanagi:2017knl, Nguyen:2017yqw], odd entanglement entropy [Tamaoka:2018ned], entanglement negativity [Kudler-Flam:2018qjo, Kusuki:2019zsp, KumarBasak:2020eia, KumarBasak:2021lwm] and widely studied at various occasions, see [BabaeiVelni:2019pkw, Jokela:2019ebz, Jeong:2019xdr, Jain:2020rbb, Basu:2022nds, Jain:2022hxl, Jain:2022csf, Jiang:2024ijx]. In [Dutta:2019gen] authors proposed a relation between reflected entropy and EWCS via the following relation
$S_{R}(A_{1},A_{2})=2E_{W}(A_{1},A_{2}).$ (3.15)
Now consider two identical parallel strip subsystems $A_{1}$ and $A_{2}$ on the boundary, each of width $l$ and separated by a distance $D$. On a constant time slice and at $x_{1}=0$, the induced metric reduces to
$ds^{2}=\frac{R^{2}}{z^{2}}\left[\frac{dz^{2}}{h(z)}+d\vec{x}^{\,2}_{d-2}\right].$ (3.16)
The EWCS between $A_{1}$ and $A_{2}$ is defined as the minimal area surface $\Gamma_{A_{1}A_{2}}$ that connects the two RT surfaces $\gamma_{A_{1}}$ and $\gamma_{A_{2}}$ bounding the entanglement wedge of $A_{1}\cup A_{2}$. Its area is given by
$E_{W}(A_{1},A_{2})=\frac{\mathcal{A}(\Gamma_{A_{1}A_{2}})}{4G_{N}^{d+1}}=\frac{R^{d-1}L^{d-2}}{4G_{N}}\int_{z_{D}}^{z_{t}}\frac{dz}{z^{d-1}\sqrt{h(z)}},$ (3.17)
where $z_{t}$ and $z_{D}$ denote the turning points corresponding to the strip width $l$ and separation $D$, respectively. These turning points are determined by eq.˜3.6. With an analogous expression for $D$ by replacing $z_{t}\to z_{D}$. Similar to the case of HEE, its analytic form is possible only in the low and high temperature limits.
Low-temperature EWCS:
In the regime $T_{f}l\ll 1$, the turning points satisfy $z_{t},z_{D}\ll z_{H}$, allowing a perturbative expansion of the lapse function $h(z)$. Substituting this into (3.17) yields
$\begin{split}E_{W}^{low}&=\frac{R^{d-1}L^{d-2}}{4G_{N}}\Bigg[K_{0}\!\left(\frac{1}{D^{d-2}}-\frac{1}{(D+2l)^{d-2}}\right)+K_{1}\,\rho\,l\!\left(\frac{4\pi T_{f}}{d}\right)^{\!d-1}\\ &\qquad\qquad\quad-\,K_{2}(1-\rho)\,l(l+D)\!\left(\frac{4\pi T_{f}}{d}\right)^{\!d}+\mathcal{O}\big((T_{f}l)^{2d-2}\big)\Bigg],\end{split}$ (3.18)
where the coefficients $K_{0},K_{1},K_{2}$ depend only on the dimension $d$ and are given explicitly in [Chakrabortty:2020ptb]. The first term corresponds to the vacuum contribution, while the $\rho$-dependent terms encode the leading and subleading backreaction effects. Notably, the correction proportional to $\rho T_{f}^{d-1}$ enhances the EWCS, indicating stronger mixed-state correlations in the presence of quark backreaction at low temperature.
High-temperature EWCS:
In the opposite limit $T_{f}l\gg 1$, the turning points approach the horizon, $z_{t},z_{D}\to z_{H}$, and the EWCS takes following the form
$\begin{split}E_{W}^{high}&=\frac{L^{d-2}R^{d-1}}{4G^{d+1}_{N}}T_{f}^{d-2}\Biggl\{-\left(\frac{2\sqrt{\pi}\Gamma\left(\frac{d}{2d-2}\right)}{\Gamma\left(\frac{1}{2d-2}\right)}\right)^{d-2}\left(\frac{1}{DT_{f}}\right)^{d-2}+\mathcal{C}\left(\frac{4\pi}{d}\right)^{d-2}\\ &\quad-\rho\left(\frac{d-2}{8\sqrt{\pi}d(d-1)}\right)\!\left(\frac{\Gamma\left(\frac{1}{2d-2}\right)}{\Gamma\left(\frac{d}{2d-2}\right)}\right)^{3}\!\left(\frac{4\pi}{d}\right)^{d-1}\!(DT_{f})\\ &\quad-(1-\rho)\left(\frac{d-2}{8\pi(d+1)}\right)\!\frac{\Gamma\left(\frac{2d}{2d-2}\right)}{\Gamma\left(\frac{1}{2}+\frac{1}{d-1}\right)}\!\left(\frac{\Gamma\left(\frac{1}{2d-2}\right)}{\Gamma\left(\frac{d}{2d-2}\right)}\right)^{3}\!\left(\frac{4\pi}{d}\right)^{d}(DT_{f})^{2}\Biggr\},\end{split}$ (3.19)
with $\mathcal{C}$ is a function of $\rho$ and one can read form [Chakrabortty:2020ptb]. In this regime, EWCS scales as $T_{f}^{d-2}$, and the backreaction enhances its magnitude monotonically. In the limit $\rho\to 0$, both (3.18) and (3.19) reduce to the AdS–Schwarzschild results, providing a nontrivial consistency check.
Reflected entropy:
In holographic theories, the reflected entropy $S_{R}(A_{1},A_{2})$ is proposed to be related to the EWCS via (3.15) [Dutta:2019gen]. This relation is supported by explicit checks in static and time-dependent AdS3/CFT2 setups as well as tensor-network constructions [Nguyen:2017yqw, Takayanagi:2017knl, Umemoto:2018jpc]. Consequently, the low and high temperature expressions for reflected entropy follow directly from (3.18) and (3.19) by an overall factor of two.
At low temperature, the leading backreaction correction $\propto\rho T_{f}^{d-1}$ enhances $S_{R}$, indicating that quark backreaction strengthens mixed state correlations beyond the vacuum contribution. At high temperature, $S_{R}$ inherits the thermal scaling $S_{R}\sim T_{f}^{d-2}$, with all backreaction dependent coefficients increasing monotonically with $\rho$. In the vanishing backreaction limit, $S_{R}$ reproduces the AdS-Schwarzschild results, confirming the robustness of the $S_{R}=2\,E_{W}$ prescription.
Although reflected entropy is geometrically determined by EWCS at the classical level, its interpretation as the entanglement entropy of a canonical purification provides additional insight beyond the purely geometric picture. In particular, while EWCS diagnoses the connectivity of the entanglement wedge, $S_{R}$ quantifies the total mixed-state entanglement shared between the subsystems, including correlations inaccessible to entanglement negativity. In the backreacted AdS black hole background, this distinction becomes important: EWCS and consequently $S_{R}$ is enhanced by backreaction at high temperature, whereas HEN exhibits a qualitatively different behavior in the low-temperature regime. This contrast highlights the complementary roles of EWCS, reflected entropy, and negativity in characterizing the entanglement structure of strongly coupled plasmas.
4 Entanglement thermodynamics
In this section, we restrict our analysis to $(3+1)$ dimensions (with $d=3$) for simplicity, although the results can be extended to higher dimensions. The metric of the deformed AdS4 black hole with a planar horizon from eq.˜2.8 takes the following form
$ds^{2}=\frac{R^{2}}{z^{2}}\left(-h(z)dt^{2}+d\vec{x}^{2}+\frac{dz^{2}}{h(z)}\right),$ (4.1)
with the blackening or lapse function expressed as
$h(z)=1-\frac{bz^{2}}{R^{2}}-\frac{2m\,z^{3}}{R^{4}}.$ (4.2)
The bulk theory is characterized by two primary parameters: the backreaction parameter $b$ and the black hole mass $m$ as described in section˜2. These parameters are related via the horizon radius $r_{H}$ through the condition $h(r_{H})=0$ yielding
$m=\frac{R^{2}\left(R^{2}-bz_{H}^{2}\right)}{2z_{H}^{3}}.$ (4.3)
The Hawking temperature for this deformed black hole can be expressed as
$T_{H}=\frac{1}{4\pi z_{H}}\left(3-\frac{bz_{H}^{2}}{R^{2}}\right).$ (4.4)
By substituting the mass relation from eq.˜4.3 into eq.˜4.2, we can express the lapse function $h(z)$ solely in terms of the horizon radius $z_{H}$ and the backreaction parameter $b$ as follows
$h(z)=1-\frac{z^{3}}{z_{H}^{3}}+\frac{bz^{2}(z-z_{H})}{R^{2}z_{H}}.$ (4.5)
This reparametrization reduces the effective parameters of the bulk theory to the backreaction $b$ and the horizon radius $z_{H}$. We then simplify the integrals for the HEE by defining $u=z/z_{t}$. Under this transformation, the expressions for the subsystem length $l$ and area $\mathcal{A}$ given in eq.˜3.6 and eq.˜3.7 respectively, leads to the following modified integral expressions
$\frac{l}{2}$ $=z_{t}\int_{0}^{1}\frac{u^{2}}{\sqrt{1-u^{4}}}\left(1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}+\frac{bz_{H}^{2}}{R^{2}}\left(\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}-\frac{u^{2}z_{t}^{2}}{z_{H}^{2}}\right)\right)^{-\frac{1}{2}}du,$ (4.6)
${\cal A}$ $=\frac{2LR^{2}}{z_{t}}\int_{0}^{1}\frac{}{u^{2}\sqrt{1-u^{4}}}\left(1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}+\frac{bz_{H}^{2}}{R^{2}}\left(\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}-\frac{u^{2}z_{t}^{2}}{z_{H}^{2}}\right)\right)^{-\frac{1}{2}}du.$ (4.7)
We utilize the subsystem length and area formulas given in eq.˜4.6 and eq.˜4.7 to investigate the thermal and backreaction effects on HEE. For a field theory dual to a deformed AdS4 spacetime, the space of states is is characterized by both $T$ (via $z_{H}$) and $b$. This is in contrast to the Schwarzschild black hole in the $AdS_{4}$ bulk scenario where $T$ is the unique scale controlling the high and low temperature limits of the entropy. In contrast, for the deformed black hole, the backreaction parameter $b$ also plays a role, as is evident from the requirement of temperature positivity in the bulk which follows from eq.˜4.4 as
$z_{H}\leq\frac{\sqrt{3}R}{\sqrt{b}}.$ (4.8)
In the relation above, the equality corresponds to the zero-temperature limit, while the inequality represents a black hole with finite temperature. According to the gauge/gravity duality, the field theory dual to a deformed AdS black hole occupies its ground state at zero temperature and enters an excited state at non-zero temperature. To derive a first law of entanglement thermodynamics, we also analyze the HEE of the dual field theory at zero temperature. This situation bears resemblance to the case of extremal and non-extremal Reissner–Nordström black holes, however a key distinction is that in the present setup the black hole mass is negative at zero temperature as discussed in section˜2, whereas it remains positive in the extremal case. We show that by defining this negative mass state as the reference ground state, the entanglement entropy satisfies a first law like behavior at low temperatures.
From the inequality in eq.˜4.8, one finds that the horizon radius $(z_{H})$ is bounded from above by a quantity inversely proportional to the backreaction parameter $b$ of the deformed black hole. Consequently, the allowed range of the horizon radius is determined by the magnitude of the backreaction. When the backreaction is small, the horizon radius $(z_{H})$ can take both large values, corresponding to the low temperature regime and small values, corresponding to the high temperature regime. In contrast, for zero-temperature deformed black holes, the equality condition in eq.˜4.8 implies that the horizon radius $(z_{H})$ is inversely related to the backreaction parameter $b$. As a result, a small backreaction leads to a large horizon radius in the zero-temperature case. In view of these observations, we examine the behavior of HEE in different regimes of the backreaction parameter for deformed AdS4 black holes. Note that in section˜3.1 the HEE was analyzed at low and high effective temperatures while keeping the dimensionless backreaction parameter $\rho$ fixed. In the present section, however we focus on the low backreaction regime and study the HEE at both low and high temperatures, allowing us to explore its dependence on both parameters and to investigate the emergence of a first law like relation. We begin by analyzing the HEE of the subsystem $A$ in the boundary field theory dual to deformed black holes at zero temperature in the small backreaction limit. We then extend the analysis to deformed black holes at non-zero temperature in the high temperature regime, again restricting to small backreaction limit.
4.1 Zero temperature
In this section, we compute the HEE of subsystem $A$ in the boundary field theory dual to a deformed $\text{AdS}_{4}$ black hole at zero temperature. By imposing the condition $T_{H}=0$ on the temperature expression in eq.˜4.4, we obtain the following constraint relating the backreaction parameter to the horizon radius as follows
$b=\frac{3R^{2}}{z_{H}^{2}}.$ (4.9)
Thus, for deformed black holes zero temperature, the small backreaction regime also corresponds to a large horizon radius $(l^{2}b\leq\frac{3R^{2}l^{2}}{z_{H}^{2}}<<1)$. With this condition, the lapse function from eq.˜4.2 simplifies
$h(z)=1-\frac{3z^{2}}{z_{H}^{2}}+\frac{2z^{3}}{z_{H}^{3}}.$ (4.10)
Substituting this form of the lapse function into eq.˜4.6 and eq.˜4.7, the integral expressions for the subsystem length $l$ and the extremal area $\mathcal{A}$ gives the following form
$\frac{l}{2}$ $=z_{t}\int_{0}^{1}\frac{u^{2}}{\sqrt{1-u^{4}}}\left(1-\frac{3u^{2}z_{t}^{2}}{z_{H}^{2}}+\frac{2u^{3}z_{t}^{3}}{z_{H}^{3}}\right)^{-\frac{1}{2}}du,$ (4.11)
${\cal A}$ $=\frac{2LR^{2}}{z_{t}}\int_{0}^{1}\frac{}{u^{2}\sqrt{1-u^{4}}}\left(1-\frac{3u^{2}z_{t}^{2}}{z_{H}^{2}}+\frac{2u^{3}z_{t}^{3}}{z_{H}^{3}}\right)^{-\frac{1}{2}}du.$ (4.12)
For large horizon radius $z_{H}$, the black hole lies deep in the bulk and is far from the extremal surface, so that $z_{t}<<z_{H}$. Consequently, we can expand $h(u)^{-1/2}$ as a power series in $z_{t}/z_{H}$ around $\frac{z_{t}}{z_{H}}=0$. Keeping terms up to the fourth order, the expansion is given by
$h(u)^{-\frac{1}{2}}\approx 1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}+\frac{3u^{2}z_{t}^{2}}{2z_{H}^{2}}.$ (4.13)
Now solving the length integral in eq.˜4.11 with the above eq.˜4.13 perturbatively in terms of $(l/z_{H})$ leads to the following expression for the turning point as
$z_{t}=\frac{2l\Gamma\left(\frac{5}{4}\right)}{\sqrt{\pi}\Gamma\left(\frac{3}{4}\right)}-\frac{16l^{3}\Gamma\left(\frac{5}{4}\right)^{5}}{\pi^{3/2}\Gamma\left(\frac{3}{4}\right)^{5}z_{H}^{2}}+\frac{8l^{4}\Gamma\left(\frac{5}{4}\right)^{5}}{\pi^{3/2}\Gamma\left(\frac{3}{4}\right)^{5}z_{H}^{3}}+\mathcal{O}(\left(\frac{z_{t}}{z_{H}}\right)^{4}).$ (4.14)
Using the same approximation for $h(u)^{-1/2}$ as in eq.˜4.13, we obtain an analytic expression for the extremal area by substituting it into eq.˜4.12 yielding
${\cal A}$ $\approx\frac{2LR^{2}}{z_{t}}\int_{0}^{1}\frac{1}{u^{2}\sqrt{1-u^{4}}}\left(1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}+\frac{3u^{2}z_{t}^{2}}{2z_{H}^{2}}\right)du$ (4.15)
$\approx\frac{2LR^{2}}{z_{t}}\left(\int_{0}^{1}\frac{1}{u^{2}\sqrt{1-u^{4}}}du-\int_{0}^{1}\frac{1}{u^{2}\sqrt{1-u^{4}}}\left(\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}-\frac{3u^{2}z_{t}^{2}}{2z_{H}^{2}}\right)du\right).$
The first term in the expression for $\mathcal{A}$ corresponds to the divergent contribution which is same as of pure AdS4. We regularize the integral by imposing a UV cutoff $\epsilon=z_{b}/z_{t}$ and incorporate a counterterm $(-2LR^{2}_{AdS})/z_{b}$ to subtract the divergence. This procedure gives the following finite value for the area as
${\cal A}^{finite}$ $=\frac{2LR^{2}}{z_{t}}\int_{\frac{z_{b}}{z_{t}}}^{1}\frac{1}{u^{2}\sqrt{1-u^{4}}}du-\frac{2LR^{2}}{z_{b}}+\frac{2LR^{2}}{z_{t}}\int_{0}^{1}\frac{1}{\sqrt{1-u^{4}}}\left(\frac{3z_{t}^{2}}{2z_{H}^{2}}-\frac{uz_{t}^{3}}{z_{H}^{3}}\right)du$ (4.16)
$=\frac{LR^{2}}{2z_{t}}\bigg[\frac{\sqrt{\pi}\Gamma(-\frac{1}{4})}{\Gamma(\frac{1}{4})}+\frac{3\sqrt{\pi}\Gamma\left(\frac{1}{4}\right)z_{t}^{2}}{2\Gamma\left(\frac{3}{4}\right)z_{H}^{2}}-\frac{\pi z_{t}^{3}}{z_{H}^{3}}\bigg].$
After substituting the turning point $z_{t}$ from eq.˜4.14 and performing a series expansion in $l/z_{H}$, the finite part of the extremal area (eq.˜4.16) leads to the following
${\cal A}^{finite}=\frac{LR^{2}}{l}\left[\frac{\pi\Gamma\left(\frac{3}{4}\right)\Gamma\left(-\frac{1}{4}\right)}{\Gamma\left(\frac{1}{4}\right)^{2}}+\frac{\pi\sqrt{2}l^{2}\Gamma\left(\frac{5}{4}\right)}{\Gamma\left(\frac{3}{4}\right)^{3}z_{H}^{2}}+\frac{16\pi\sqrt{2}l^{3}\Gamma\left(\frac{5}{4}\right)}{\Gamma\left(-\frac{1}{4}\right)^{3}z_{H}^{3}}+\mathcal{O}(\frac{l}{z_{H}})^{4}\right].$ (4.17)
The renormalized HEE of the subsystem $A$ in the boundary theory dual to the deformed black hole at zero temperature in the small backreaction regime can be obtained using the RT formula as
$S_{A}^{finite}=\frac{1}{4G_{N}^{4}}\frac{LR^{2}}{l}\left[\frac{\pi\Gamma\left(\frac{3}{4}\right)\Gamma\left(-\frac{1}{4}\right)}{\Gamma\left(\frac{1}{4}\right)^{2}}+\frac{\pi\sqrt{2}\Gamma\left(\frac{5}{4}\right)b\,l^{2}}{3R^{2}_{AdS}\Gamma\left(\frac{3}{4}\right)^{3}}+\frac{16\pi\sqrt{2}l^{3}\Gamma\left(\frac{5}{4}\right)}{\Gamma\left(-\frac{1}{4}\right)^{3}z_{H}^{3}}+\mathcal{O}(\frac{l}{z_{H}})^{4}\right].$ (4.18)
Therefore, in the zero-temperature case, the HEE receives corrections that are linear and quadratic in $l$ and $l^{2}$. Notably, the emergence of a term linear in $l$ distinguishes this geometry from the extremal RN black hole case [Chaturvedi:2016kbk] where it is absent. By substituting the zero temperature mass $m_{0}=-R^{4}/z_{H}^{3}$ obtained from eq.˜4.3 and eq.˜4.9 into above eq.˜4.18, the renormalized HEE can be expressed in the following compact form
$S_{A}^{finite}\approx S_{A}^{AdS}+k_{1}bLl+k_{2}m_{0}Ll^{2},$ (4.19)
where $S_{A}^{AdS}$ denotes the HEE for pure AdS4 and constants $k_{1,2}$ are given by
$k_{1}=\frac{1}{4G_{N}^{4}}\frac{\pi\sqrt{2}\Gamma\left(\frac{5}{4}\right)}{3\Gamma\left(\frac{3}{4}\right)^{3}},\quad k_{2}=\frac{1}{4G_{N}^{4}R^{2}}\frac{\Gamma\left(\frac{5}{4}\right)^{2}}{\Gamma\left(\frac{3}{4}\right)^{2}}.$ (4.20)
We see that in the small backreaction regime, the dominant contribution to the HEE of subsystem $A$ comes from the pure AdS4 geometry. The subleading corrections in the eq.˜4.19 play a crucial role in subsequent sections where they form the basis for establishing a first law like relation for the HEE.
4.2 Low temperature regime
We now turn to the analysis of the subsystem $A$ in the boundary field theory dual to a deformed $AdS_{4}$ black hole with small backreaction at low temperature. As follows from eq.˜4.8, when both the temperature and the backreaction are small, the horizon radius $z_{H}$ becomes large. In this limit, one has $bz_{H}^{2}/R^{2}\sim 1$ and $z_{H}$ sufficiently large such that $z_{t}\ll z_{H}$. Under these conditions, the lapse function $h(u)$ takes the form
$h(u)=1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}+\frac{bz_{H}^{2}}{R^{2}}\left(\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}-\frac{u^{2}z_{t}^{2}}{z_{H}^{2}}\right).$ (4.21)
By introducing the dimensionless parameter $\alpha=bz_{H}^{2}/R^{2}$, we perform a Taylor expansion of $h(u)^{-1/2}$ around the $\frac{z_{t}}{z_{H}}=0$. Neglecting higher order terms, we obtain the following approximated expression
$h(u)^{-\frac{1}{2}}\approx 1+\frac{\alpha u^{2}z_{t}^{2}}{2z_{H}^{2}}+\frac{\left(1-\alpha\right)u^{3}z_{t}^{3}}{2z_{H}^{3}}.$ (4.22)
On utilizing the above expression in eq.˜4.6 for subsystem length $l$, we obtain
$\frac{l}{2}\approx z_{t}\int_{0}^{1}\frac{u^{2}}{\sqrt{1-u^{4}}}\left(1+\frac{\alpha u^{2}z_{t}^{2}}{2z_{H}^{2}}+\frac{\left(1-\alpha\right)u^{3}z_{t}^{3}}{2z_{H}^{3}}\right)du.$ (4.23)
By inverting the above eq.˜4.23, we express $z_{t}$ as a function of the subsystem length $l$. Solving this relation perturbatively in the small parameter $(l/z_{H})$ yields the following expression for the turning point as
$z_{t}=\frac{2l\Gamma\left(\frac{5}{4}\right)}{\sqrt{\pi}\Gamma\left(\frac{3}{4}\right)}-\frac{27\alpha l^{3}\Gamma\left(\frac{5}{4}\right)^{5}}{16\pi^{3/2}\Gamma\left(\frac{3}{4}\right)\Gamma\left(\frac{7}{4}\right)^{4}z_{H}^{2}}+\frac{4(\alpha-1)l^{4}\Gamma\left(\frac{5}{4}\right)^{5}}{\pi^{3/2}\Gamma\left(\frac{3}{4}\right)^{5}z_{H}^{3}}+\mathcal{O}[(\frac{l}{z_{H}})^{4}].$ (4.24)
The extremal area is obtained by applying the approximation in eq.˜4.22 to the integral defined in eq.˜4.7 as follows
${\cal A}\approx\frac{2LR^{2}}{z_{t}}\int_{0}^{1}\frac{1}{u^{2}\sqrt{1-u^{4}}}\left(1+\frac{\alpha u^{2}z_{t}^{2}}{2z_{H}^{2}}+\frac{\left(1-\alpha\right)u^{3}z_{t}^{3}}{2z_{H}^{3}}\right)du.$ (4.25)
The expression for the extremal area in eq.˜4.25 contains a divergent term identical to that of the pure AdS4 vacuum. By applying the same regularization procedure used for the zero temperature case, we regularize the integral to isolate the finite contribution. The resulting finite part of the area is then given by
${\cal A}^{finite}\approx LR^{2}\left[\frac{\sqrt{\pi}\Gamma(-\frac{1}{4})}{2\Gamma(\frac{1}{4})z_{t}}+\frac{\sqrt{\pi}\alpha\Gamma\left(\frac{5}{4}\right)z_{t}}{\Gamma\left(\frac{3}{4}\right)z_{H}^{2}}+\frac{\pi(\alpha-1)z_{t}^{2}}{4z_{H}^{3}}\right].$ (4.26)
Now substituting $z_{t}$ from eq.˜4.24 into eq.˜4.26 and truncating the expansion at fourth order in $(l/z_{H})$, the finite contribution to the extremal area becomes
${\cal A}^{finite}=\frac{LR^{2}}{l}\left[\frac{\pi\Gamma\left(\frac{3}{4}\right)\Gamma\left(-\frac{1}{4}\right)}{\Gamma\left(\frac{1}{4}\right)^{2}}+\frac{9\pi^{2}\alpha l^{2}}{128\Gamma\left(\frac{3}{4}\right)\Gamma\left(\frac{7}{4}\right)^{3}z_{H}^{2}}-\frac{8(\alpha-1)l^{3}\Gamma\left(\frac{5}{4}\right)^{2}}{\Gamma\left(-\frac{1}{4}\right)^{2}z_{H}^{3}}+\mathcal{O}[(\frac{l}{z_{H}})^{4}\right].$ (4.27)
The holographic renormalized entanglement entropy of the subsystem $A$ for small backreaction at low temperature regime may now be obtained using the RT formula as follows
$S_{A}^{finite}=\frac{LR^{2}}{4G_{N}^{4}\,l}\left[\frac{\pi\Gamma\left(\frac{3}{4}\right)\Gamma\left(-\frac{1}{4}\right)}{\Gamma\left(\frac{1}{4}\right)^{2}}+\frac{\sqrt{2}\pi\,\Gamma\left(\frac{5}{4}\right)b\,l^{2}}{3R^{2}\Gamma\left(\frac{3}{4}\right)^{3}}+\frac{8(1-\alpha)l^{3}\Gamma\left(\frac{5}{4}\right)^{2}}{\Gamma\left(-\frac{1}{4}\right)^{2}z_{H}^{3}}+\mathcal{O}[(\frac{l}{z_{H}})^{4}]\right].$ (4.28)
On using the expression for mass from eq.˜4.3, the above expression can be written in compact form as
$S_{A}^{finite}\approx S_{A}^{AdS}+k_{1}bLl+k_{2}mLl^{2},$ (4.29)
where the constants $k_{1}$ and $k_{2}$ are same given in eq.˜4.20. We see that when backreaction and temperature both are small, the dominant contribution to the HEE for deformed black holes at non-zero temperature remains the pure AdS4 similar to the case of zero temperature black holes.
4.3 First law of entanglement thermodynamics
We now examine entanglement thermodynamics within the small backreaction regime. According to the AdS/CFT correspondence, the deformed AdS black hole at zero temperature is identified as the holographic dual to the ground state of the boundary field theory. In contrast, the deformed black hole at finite temperature corresponds to an excited state. The change in HEE $\Delta S_{A}$ between the excited and ground states leads to a first law of entanglement thermodynamics as follows
$\Delta E_{A}=T_{\mathrm{ent}}\Delta S_{A},$ (4.30)
where $T_{\mathrm{ent}}$ is known as the entanglement temperature and the change in HEE is computed by taking the difference between eq.˜4.29 and eq.˜4.19 yielding
$\Delta S_{A}=S_{A}-S_{A}^{T=0}=(m-m_{0})k_{2}Ll^{2}.$ (4.31)
Furthermore, the energy increment associated with subsystem $A$ is defined as
$\Delta E_{A}=\int_{A}dx^{2}\ T_{tt}^{T_{H}\neq 0}-\int_{A}dx^{2}\ T_{tt}^{T_{H}=0}=\frac{Ll}{4\pi G_{N}^{4}R^{2}}(m-m_{0}),$ (4.32)
where $T_{tt}$ denotes the time component of the boundary stress-energy tensor and can be obtained using the holographic prescription of [Balasubramanian:1999re]. Now substituting the results from eq.˜4.31 and eq.˜4.32 into the first law like relation eq.˜4.30, the entanglement temperature is found to be
$T_{\mathrm{ent}}=\frac{\Delta E_{A}}{\Delta S_{A}}=\frac{1}{l}\frac{16\,\Gamma\left(\frac{3}{4}\right)^{2}}{\pi\,\Gamma\left(\frac{1}{4}\right)^{2}}.$ (4.33)
We observe that the entanglement temperature $T_{\mathrm{ent}}$ is proportional to the inverse of the subsystem size. This universal scaling behavior is consistent with the entanglement thermodynamics of the asymptotically AdS geometries as discussed in [Bhattacharya:2012mi]. This suggests that the small backreaction does not alter the fundamental scaling of the first law and the HEE continues to satisfy a well defined thermodynamic relation in the perturbative regime.
4.4 High temperature regime
In this section, we analyze the high temperature behavior of the HEE for a subsystem $A$ when the bulk deformed black hole has a small backreaction and is in the high-temperature regime. From eq.˜4.8, it follows that in the limit of high temperature ($l/z_{H}>>1$) and small backreaction, the parameter $bz_{H}^{2}/3R^{2}$ becomes much smaller than unity i.e $\frac{bz_{H}^{2}}{3R^{2}}\ll 1$. By introducing the small parameter $\delta=bz_{H}^{2}/3R^{2}$, we can Taylor expand the lapse function $h(u)^{-1/2}$ around $\delta=0$ yielding
$h(u)^{-\frac{1}{2}}\approx\frac{1}{\sqrt{1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}}}+\frac{3\frac{u^{2}z_{t}^{2}}{z_{H}^{2}}\delta\left(1-\frac{uz_{t}}{z_{H}}\right)}{2\left(1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}\right){}^{3/2}}.$ (4.34)
By employing the approximation for $h(u)$ given in eq.˜4.34, the subsystem length $l$ and the extremal surface area $\mathcal{A}$ are evaluated using eq.˜4.6 and eq.˜4.7 respectively as follows
$l$ $=2z_{t}\int_{0}^{1}\frac{u^{2}}{\sqrt{1-u^{4}}}\left(\frac{1}{\sqrt{1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}}}+\frac{3\delta u^{2}z_{t}^{2}\left(1-\frac{uz_{t}}{z_{H}}\right)}{2z_{H}^{2}\left(1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}\right){}^{3/2}}\right)\,du$ (4.35)
$\mathcal{A}$ $=\frac{2LR^{2}}{z_{t}}\int_{0}^{1}\frac{1}{u^{2}\sqrt{1-u^{4}}}\left(\frac{1}{\sqrt{1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}}}+\frac{3\delta u^{2}z_{t}^{2}\left(1-\frac{uz_{t}}{z_{H}}\right)}{2z_{H}^{2}\left(1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}\right){}^{3/2}}\right)\,du$ (4.36)
The integrals above do not admit simple closed forms. To proceed, we utilize the following identities:
$\frac{1}{\sqrt{1-x}}=\sum_{n=0}^{\infty}\frac{\Gamma(n+\frac{1}{2})}{\sqrt{\pi}\Gamma(n+1)}x^{n},\frac{1}{{(1-x)}^{\frac{3}{2}}}=\sum_{n=0}^{\infty}\frac{2\Gamma(n+\frac{3}{2})}{\sqrt{\pi}\Gamma(n+1)}x^{n},\int_{0}^{1}\frac{u^{m}}{\sqrt{1-u^{4}}}du=\frac{\sqrt{\pi}}{4}\frac{\Gamma(\frac{m+1}{4})}{\Gamma(\frac{m+3}{4})},$ (4.37)
to express the integrals into a more tractable form. Consequently, the expression for the subsystem length given in eq.˜4.35 can be rewritten as
$\frac{l}{2z_{t}}=$ $\sum_{n=0}^{\infty}\frac{\Gamma(n+\frac{1}{2})}{\sqrt{\pi}\Gamma(n+1)}\left(\frac{z_{t}}{z_{H}}\right)^{3n}\int_{0}^{1}\frac{u^{3n+2}}{\sqrt{1-u^{4}}}+\delta\sum_{n=0}^{\infty}\frac{3\Gamma(n+\frac{3}{2})}{\sqrt{\pi}\Gamma(n+1)}\left(\frac{z_{t}}{z_{H}}\right)^{3n+2}\int_{0}^{1}\frac{u^{3n+4}(1-u\frac{z_{t}}{z_{H}})}{\sqrt{1-u^{4}}}$ (4.38)
$\frac{l}{z_{t}}=$ $\sum_{n=0}^{\infty}\frac{\Gamma(n+\frac{1}{2})}{2\Gamma(n+1)}\frac{\Gamma(\frac{3n+3}{4})}{\Gamma(\frac{3n+5}{4})}\left(\frac{z_{t}}{z_{H}}\right)^{3n}+\frac{3\delta}{2}\sum_{n=0}^{\infty}\frac{\Gamma(n+\frac{3}{2})}{\Gamma(n+1)}\frac{\Gamma(\frac{3n+5}{4})}{\Gamma(\frac{3n+7}{4})}\left(\frac{z_{t}}{z_{H}}\right)^{3n+2}$
$\quad-\frac{3\delta}{2}\sum_{n=0}^{\infty}\frac{\Gamma(n+\frac{3}{2})}{\Gamma(n+1)}\frac{\Gamma(\frac{3n+6}{4})}{\Gamma(\frac{3n+8}{4})}\left(\frac{z_{t}}{z_{H}}\right)^{3n+3}.$
Let us now examine the asymptotic behavior of the terms in the above expression to identify potential divergences. For large $n$, the first series scales as $\frac{1}{\sqrt{3}\,n}(z_{t}/z_{H})^{3n}$, while the remaining two series behave as $\frac{2}{\sqrt{3}}(z_{t}/z_{H})^{3n}$. Upon isolating these terms, it becomes evident that the divergences in the latter two series cancel one another, leading to the following simplified expression as
$\frac{l}{z_{t}}=$ $\frac{\sqrt{\pi}}{2}\frac{\Gamma(\frac{3}{4})}{\Gamma(\frac{5}{4})}+\sum_{n=1}^{\infty}\left(\frac{\Gamma(n+\frac{1}{2})}{2\Gamma(n+1)}\frac{\Gamma(\frac{3n+3}{4})}{\Gamma(\frac{3n+5}{4})}-\frac{1}{\sqrt{3}n}\right)\left(\frac{z_{t}}{z_{H}}\right)^{3n}$ (4.39)
$+\frac{3\delta}{2}\sum_{n=0}^{\infty}\left(\frac{\Gamma(n+\frac{3}{2})}{\Gamma(n+1)}\frac{\Gamma(\frac{3n+5}{4})}{\Gamma(\frac{3n+7}{4})}-\frac{2}{\sqrt{3}}\right)\left(\frac{z_{t}}{z_{H}}\right)^{3n+2}-\frac{3\delta}{2}\sum_{n=0}^{\infty}\left(\frac{\Gamma(n+\frac{3}{2})}{\Gamma(n+1)}\frac{\Gamma(\frac{3n+6}{4})}{\Gamma(\frac{3n+8}{4})}-\frac{2}{\sqrt{3}}\right)\left(\frac{z_{t}}{z_{H}}\right)^{3n+3}$
$-\frac{1}{\sqrt{3}}\log\left[1-\left(\frac{z_{t}}{z_{H}}\right)^{3}\right]+\sqrt{3}\delta\frac{(\frac{z_{t}}{z_{H}})^{2}}{(1+\frac{z_{t}}{z_{H}}+(\frac{z_{t}}{z_{H}})^{2})}.$
As described in[Hubeny:2012ry], for a static asymptotically AdS black hole, the extremal surface cannot penetrate the event horizon, necessitating that $z_{t}<z_{H}$. In the high temperature limit, however, the turning point approaches the horizon such that $z_{t}\to z_{H}$. By introducing a small positive parameter $\epsilon\ll 1$, we can parameterize this proximity as $z_{t}=(1-\epsilon)z_{H}$. Taking the limit $\epsilon\to 0$ in eq.˜4.39, the series converges to the following compact leading order expression as follows
$\frac{l}{z_{H}}=-\frac{1}{\sqrt{3}}\log[3\epsilon]+C_{1}+\delta C_{2}+\mathcal{O}[\epsilon],$ (4.40)
where the coefficients $c_{1}$ and $c_{2}$ are given by
$C_{1}=$ $\frac{\sqrt{\pi}}{2}\frac{\Gamma(\frac{3}{4})}{\Gamma(\frac{5}{4})}+\sum_{n=1}^{\infty}\bigg(\frac{\Gamma(n+\frac{1}{2})}{2\Gamma(n+1)}\frac{\Gamma(\frac{3n+3}{4})}{\Gamma(\frac{3n+5}{4})}-\frac{1}{\sqrt{3}n}\bigg)$ (4.41)
$C_{2}=$ $\frac{1}{\sqrt{3}}+\frac{3}{2}\sum_{n=0}^{\infty}\bigg(\frac{\Gamma(n+\frac{3}{2})}{\Gamma(n+1)}\frac{\Gamma(\frac{3n+5}{4})}{\Gamma(\frac{3n+7}{4})}-\frac{2}{\sqrt{3}}\bigg)-\frac{3}{2}\sum_{n=0}^{\infty}\bigg(\frac{\Gamma(n+\frac{3}{2})}{\Gamma(n+1)}\frac{\Gamma(\frac{3n+6}{4})}{\Gamma(\frac{3n+8}{4})}-\frac{2}{\sqrt{3}}\bigg).$
In evaluating the area integral given in eq.˜4.36, divergent part is confined to the first term. Following [Fischler:2012ca], we isolate the divergence and express the remaining finite contribution as
${\cal A}^{finite}=\frac{2LR^{2}}{z_{t}}\int_{\frac{z_{b}}{r_{t}}}^{1}\frac{1}{u^{2}\sqrt{1-u^{4}}}\left(\frac{1}{\sqrt{1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}}}+\frac{3\delta u^{2}z_{t}^{2}\left(1-\frac{uz_{t}}{z_{H}}\right)}{2z_{H}^{2}\left(1-\frac{u^{3}z_{t}^{3}}{z_{H}^{3}}\right){}^{3/2}}\right)\,du-\frac{2L}{z_{b}}.$ (4.42)
Once again using the identities given in eq.˜4.37, the above integral reduces to following form
${\cal A}^{finite}$ $=\frac{LR^{2}}{z_{t}}\bigg[\sqrt{\pi}\frac{\Gamma(-\frac{1}{4})}{2\Gamma(\frac{1}{4})}+\sum_{n=1}^{\infty}\frac{\Gamma(n+\frac{1}{2})}{2\Gamma(n+1)}\frac{\Gamma(\frac{3n-1}{4})}{\Gamma(\frac{3n+1}{4})}\left(\frac{z_{t}}{z_{H}}\right)^{3n}$
$+\frac{3\delta}{2}\sum_{n=0}^{\infty}\frac{\Gamma(n+\frac{3}{2})}{\Gamma(n+1)}\frac{\Gamma(\frac{3n+1}{4})}{\Gamma(\frac{3n+3}{4})}\left(\frac{z_{t}}{z_{H}}\right)^{3n+2}-\frac{3\delta}{2}\sum_{n=0}^{\infty}\frac{\Gamma(n+\frac{3}{2})}{\Gamma(n+1)}\frac{\Gamma(\frac{3n+2}{4})}{\Gamma(\frac{3n+4}{4})}\left(\frac{z_{t}}{z_{H}}\right)^{3n+3}\bigg].$
After using the identity $\Gamma(n+1)=n\Gamma(n)$ in the above equation leads to the following
${\cal A}^{finite}$ $=\frac{LR^{2}}{z_{t}}\bigg[\sqrt{\pi}\frac{\Gamma(-\frac{1}{4})}{2\Gamma(\frac{1}{4})}+\sum_{n=1}^{\infty}\bigg(1+\frac{2}{3n-1}\bigg)\frac{\Gamma(n+\frac{1}{2})}{2\Gamma(n+1)}\frac{\Gamma(\frac{3n+3}{4})}{\Gamma(\frac{3n+5}{4})}\left(\frac{z_{t}}{z_{H}}\right)^{3n}$ (4.43)
$+\frac{3\delta}{2}\sum_{n=0}^{\infty}\bigg(1+\frac{2}{3n+1}\bigg)\frac{\Gamma(n+\frac{3}{2})}{\Gamma(n+1)}\frac{\Gamma(\frac{3n+5}{4})}{\Gamma(\frac{3n+7}{4})}\left(\frac{z_{t}}{z_{H}}\right)^{3n+2}$
$-\frac{3\delta}{2}\sum_{n=0}^{\infty}\bigg(1+\frac{2}{3n+2}\bigg)\frac{\Gamma(n+\frac{3}{2})}{\Gamma(n+1)}\frac{\Gamma(\frac{3n+6}{4})}{\Gamma(\frac{3n+8}{4})}\left(\frac{z_{t}}{z_{H}}\right)^{3n+3}\bigg].$
We can now use eq.˜4.38 in above eq.˜4.43 to simplify the extremal area as
${\cal A}^{finite}$ $=\frac{LR^{2}}{z_{t}}\bigg[\sqrt{\pi}\frac{\Gamma(-\frac{1}{4})}{2\Gamma(\frac{1}{4})}+\frac{l}{z_{t}}+\sum_{n=1}^{\infty}\bigg(\frac{1}{3n-1}\bigg)\frac{\Gamma(n+\frac{1}{2})}{\Gamma(n+1)}\frac{\Gamma(\frac{3n+3}{4})}{\Gamma(\frac{3n+5}{4})}\left(\frac{z_{t}}{z_{H}}\right)^{3n}$ (4.44)
$+3\delta\sum_{n=0}^{\infty}\bigg(\frac{1}{3n+1}\bigg)\frac{\Gamma(n+\frac{3}{2})}{\Gamma(n+1)}\frac{\Gamma(\frac{3n+5}{4})}{\Gamma(\frac{3n+7}{4})}\left(\frac{z_{t}}{z_{H}}\right)^{3n+2}$
$-3\delta\sum_{n=0}^{\infty}\bigg(\frac{1}{3n+2}\bigg)\frac{\Gamma(n+\frac{3}{2})}{\Gamma(n+1)}\frac{\Gamma(\frac{3n+6}{4})}{\Gamma(\frac{3n+8}{4})}\left(\frac{z_{t}}{z_{H}}\right)^{3n+3}\bigg].$
The first series in the above equation goes as $\frac{2}{3\sqrt{3}n^{2}}(z_{t}/z_{H})^{3n}$ while the latter two series scale as $\frac{2}{3\sqrt{3}n}(z_{t}/z_{H})^{3n}$ for large $n$. Given that these series diverge at $\mathcal{O}(\epsilon)$ and beyond, we separate the divergent parts to obtain the following
${\cal A}^{finite}$ $=\frac{LR^{2}}{z_{t}}\Bigg[\sqrt{\pi}\frac{\Gamma(-\frac{1}{4})}{2\Gamma(\frac{1}{4})}+\frac{l}{z_{t}}+\sum_{n=1}^{\infty}\bigg(\frac{1}{3n-1}\frac{\Gamma(n+\frac{1}{2})}{\Gamma(n+1)}\frac{\Gamma(\frac{3n+3}{4})}{\Gamma(\frac{3n+5}{4})}-\frac{2}{3\sqrt{3}n^{2}}\bigg)\left(\frac{z_{t}}{z_{H}}\right)^{3n}$ (4.45)
$+3\delta\sum_{n=1}^{\infty}\bigg(\frac{1}{3n+1}\frac{\Gamma(n+\frac{3}{2})}{\Gamma(n+1)}\frac{\Gamma(\frac{3n+5}{4})}{\Gamma(\frac{3n+7}{4})}-\frac{2}{3\sqrt{3}n}\bigg)\left(\frac{z_{t}}{z_{H}}\right)^{3n+2}$
$-3\delta\sum_{n=1}^{\infty}\bigg(\frac{1}{3n+2}\frac{\Gamma(n+\frac{3}{2})}{\Gamma(n+1)}\frac{\Gamma(\frac{3n+6}{4})}{\Gamma(\frac{3n+8}{4})}-\frac{2}{3\sqrt{3}n}\bigg)\left(\frac{z_{t}}{z_{H}}\right)^{3n+3}$
$+3\delta\bigg(\frac{\Gamma(\frac{3}{2})\Gamma(\frac{5}{4})}{\Gamma(\frac{7}{4})}\left(\frac{z_{t}}{z_{H}}\right)^{2}-\frac{(\Gamma(\frac{3}{2}))^{2}}{2}\left(\frac{z_{t}}{z_{H}}\right)^{3}\bigg)+\frac{2}{3\sqrt{3}}\mathrm{Li}_{2}\left[\left(\frac{z_{t}}{z_{H}}\right)^{3}\right]$
$-\frac{2\delta}{\sqrt{3}}(1-\frac{z_{t}}{z_{H}})\left(\frac{z_{t}}{z_{H}}\right)^{2}\log\left[1-\left(\frac{z_{t}}{z_{H}}\right)^{3}\right]\Bigg].$
Upon substituting $z_{t}=z_{H}(1-\epsilon)$ into the above expression and performing an expansion to order $\mathcal{O}(\epsilon)$ allow us to recast the result in the compact form as
${\cal A}^{finite}=\frac{LR^{2}l}{z_{H}^{2}}+\frac{LR^{2}}{z_{H}}\Bigg[A_{0}+\delta D_{0}+\epsilon\Big(A_{1}+2\frac{l}{z_{H}}+\delta D_{1}+\frac{2}{\sqrt{3}}(1-\delta)\ln\epsilon\Big)+\mathcal{O}(\epsilon^{2})\Bigg],$ (4.46)
where the constants $A_{0,1},D_{0,1}$ are detailed in the appendix A. Now using the RT formula, the HEE in the limit of high temperature and small backreaction can be expressed as follows
$S_{A}^{finite}=Ll\,s+\frac{LR^{2}}{4G_{N}^{4}z_{H}}\Bigg[A_{0}+\delta D_{0}+\epsilon\Big(A_{1}+2\frac{l}{z_{H}}+\delta D_{1}+\frac{2}{\sqrt{3}}(1-\delta)\ln\epsilon\Big)+\mathcal{O}(\epsilon^{2})\Bigg].$ (4.47)
Here the term $s=\frac{R^{2}}{4G_{N}^{4}z_{H}^{2}}$ corresponds to the entropy density of the black hole as can be seen from from the eq.˜2.13. In eq.˜4.47, the first term follows a volume law scaling (proportional to the subsystem area in the $(2+1)$ dimensional dual field theory) and grows with temperature indicating that it represents the thermal part of the entropy. The subleading terms scale with the length of the entangling surface separating subsystem $A$ from its complement and thus represents entanglement contributions at high temperatures. Notably, the $\epsilon$-corrections exhibit an exponential decay with temperature analogous to the Schwarzschild-AdS case although they are modified by a small $\delta$ term arising from the backreaction.
Holographic Entanglement negativity
Entanglement negativity is a well established quantum information theoretic measure that captures the distillable quantum correlations in mixed states, making it particularly suitable for thermal or finite-density systems where entanglement entropy alone fails to isolate genuine quantum correlations. Defined via the partial transpose of the reduced density matrix, negativity provides a computable upper bound on the distillable entanglement. Within the AdS/CFT framework, a holographic prescription for entanglement negativity in higher dimensions has been proposed, wherein suitable linear combinations of Ryu–Takayanagi surfaces encode the negativity of boundary mixed states [Chaturvedi:2016rft, Jain:2017xsu, Jain:2018bai, KumarBasak:2020viv, Mondal:2021kzj, Afrasiar:2021hld]. This construction allows one to extend the study of mixed state entanglement to strongly coupled holographic plasmas, and in what follows we analyze its behavior in the deformed AdS black hole background to compare with the trends observed for HEN.
5 HEN for adjacent subsystems
Let us consider two adjacent rectangular strip subsystems $A_{1}$ and $A_{2}$ with respective lengths $l_{1}$ and $l_{2}$ described by
$x\equiv x^{1}=\biggl[-\frac{l_{j}}{2},\frac{l_{j}}{2}\biggr],\quad x^{i}=\biggl[-\frac{L}{2},\frac{L}{2}\biggr];\quad j=1,2,$ (5.1)
where i=2,3,,d1i=2,3,...,d-1 and $L\to\infty$. The holographic prescription then requires an appropriate combination of the entanglement entropies of the individual regions and their union [Jain:2017aqk, Jain:2017xsu]. This construction effectively isolates the quantum part of correlations by eliminating purely thermal contributions. In what follows, we will evaluate the holographic negativity for the adjacent strip configuration in the backreacted AdS black hole background and examine its dependence on quark density $\rho$, with particular focus on the low and high temperature limits, in order to contrast its behavior with that of the entanglement entropy discussed previously. Following the proposal given in [Jain:2017xsu], HEN for two adjacent subsystems is given by
$\mathcal{E}=\frac{3}{16G_{N}}(\mathcal{A}_{1}+\mathcal{A}_{2}-\mathcal{A}_{3}),$ (5.2)
Here, $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, and $\mathcal{A}_{3}=\mathcal{A}_{1}+\mathcal{A}_{2}$ represent the areas of the corresponding RT surfaces, obtained using the prescription reviewed in section˜3. To express the entanglement negativity in terms of boundary parameters, one needs to determine the turning point $z_{t}$ as a function of the strip width $l$ by solving (3.6). However, as discussed in section˜3, this relation cannot be solved exactly for arbitrary $z_{t}$ and admits analytic solutions only in the low and high temperature regimes. Therefore, we evaluate the HEN in these two limiting cases using the corresponding results for the HEE.
5.1 Low temperature limit
The low temperature regime corresponds to the condition $z_{t}\ll z_{h}$. In terms of the boundary parameter, this limit translates to $Tl\ll 1$. By solving eq.˜3.6 perturbatively for $z_{t}$ in powers of $l$, one can obtain the corresponding expressions for the area functions $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, and $\mathcal{A}_{3}$ in this regime, as outlined in [Chakrabortty:2020ptb]. The resulting expression for the area is given by
$\begin{split}\mathcal{A}_{j}=\frac{2}{d-2}\frac{L^{d-2}R^{d-1}}{\epsilon^{d-2}}&+R^{d-1}\left(\frac{L}{l_{j}}\right)^{d-2}S_{0}\Biggl(1+\rho S_{2}\left(\frac{4\pi T_{f}l}{d}\right)^{d-1}\\ &+(1-\rho)S_{1}\left(\frac{4\pi T_{f}l_{j}}{d}\right)^{d}+\mathcal{O}\left((T_{f}l_{j})^{2d-2}\right)\Biggr).\end{split}$ (5.3)
Here, the index $j=1,2,3$ labels the subsystems $A_{1}$, $A_{2}$, and $A_{3}$, corresponding to their respective strip widths. In eq.˜5.3, the first term with $\epsilon$ denotes the divergent contribution, while $S_{0}$, $S_{1}$, and $S_{2}$ are constants depending only on the spacetime dimension $d$ and are given in eq.˜3.10.
Using (5.2) together with the low-temperature expansion eq.˜5.3 for the area functionals, the HEN for two adjacent strip intervals in the deformed AdS black hole background takes the form
$\begin{split}\mathcal{E}&=\frac{3}{8G_{N}(d-2)}\frac{L^{d-2}R^{d-1}}{\epsilon^{d-2}}\quad+\frac{3R^{d-1}L^{d-2}}{8G_{N}}S_{0}\Biggl(l_{1}^{2-d}+l_{2}^{2-d}-(l_{1}+l_{2})^{2-d}\\ &\qquad\qquad\qquad\qquad\qquad\qquad-\;2S_{1}(1-\rho)\,l_{1}l_{2}\Big(\tfrac{4\pi T_{f}}{d}\Big)^{\!d}+\mathcal{O}\!\big((T_{f}l_{j})^{2d-2}\big)\Biggr).\end{split}$ (5.4)
From the above expression eq.˜5.4, it follows that :
  • The first term is the standard ultraviolet divergence arising from short-distance correlations near the entangling surface. The leading finite contribution (independent of $T_{f}$) is the familiar vacuum-like short-distance scaling for adjacent strip geometries and encodes the purely geometric contribution to negativity.
  • The first nontrivial thermal/backreaction correction appears at order $T_{f}^{\,d}$ and is multiplied by the factor $(1-\rho)$. In low temperature regime $T_{f}^{\,d}$ term is the leading thermal correction; higher-order terms scale as $(T_{f}l_{j})^{2d-2}$ and are subleading for sufficiently small $T_{f}l_{j}$.
  • The correction proportional to $(1-\rho)T_{f}^{\,d}$ reduces the finite part of $\mathcal{E}$. For fixed $T_{f}$ and interval sizes the negativity therefore decreases as the backreaction increases toward unity, i.e. the magnitude of this negative correction is proportional to $(1-\rho)$ and vanishes as $\rho\to 1$. Physically, this indicates that adding heavy-flavour backreaction tends to enhance thermal/classical correlations which compete with and reduce the purely quantum correlations measured by negativity in the low-temperature regime.
  • This behavior of negativity at low $T_{f}$ is consistent with the HEE analysis where backreaction-generated terms modify the finite part of the entropy and tend to push the system toward thermal-dominated behaviour; negativity—being a quantifier of distillable entanglement-like correlations—responds by decreasing when thermal contributions grow.
The low-temperature adjacent-interval result (5.4) demonstrates that, in the near-vacuum regime, heavy-quark backreaction produces a leading thermal correction $\propto(1-\rho)T_{f}^{d}$ which lowers the HEN. This is consistent with the qualitative expectation that additional matter content increases thermal correlations and thereby diminishes quantum entanglement as measured by negativity.
5.2 High temperature limit
[…]
$\mathcal{E}\;=\;\frac{3\,L^{d-2}R^{d-1}}{8G_{N}}\,\biggl(\frac{4\pi T_{f}}{d}\biggr)^{\!d-2}\,\Big[\,\widetilde{S}_{1}(\rho,d)+\widetilde{S}_{2}(\rho,d)-\widetilde{S}_{12}(\rho,d)\,\Big]\,$, (5.5)
where the dimensionless functions $\widetilde{S}_{1},\widetilde{S}_{2},\widetilde{S}_{12}$ are defined in eq.˜3.14 and encode the backreaction dependence; numerically (and as established in [Chakrabortty:2020ptb]) each $\widetilde{S}(\rho,d)$ is a monotonically increasing function of $\rho$. From the above result, it is evident that
  • The combination of $\widetilde{S}$’s in square brackets is finite and captures all nontrivial dependence on the backreaction parameter $\rho$ and the spacetime dimension $d$. Divergent pieces present in the separate area integrals cancel in the prescribed linear combination, yielding the well-defined finite coefficient multiplying the thermal scaling.
  • Since each $\widetilde{S}(\rho,d)$ grows with $\rho$ as noticed in [Chakrabortty:2020ptb], the combination in (5.5) under squre bracket typically increases with $\rho$. Therefore, at high effective temperature the net effect of turning on quark backreaction is an increase of the finite part of $\mathcal{E}$.
  • Physically, this enhancement can be understood as a modification of the near-horizon geometry by the backreacting matter: the geometric measure that feeds into the negativity receives larger near-horizon contributions when $\rho$ is increased, which shows up as larger values of the $\widetilde{S}$ coefficients.
At high effective temperature the adjacent-interval HEN becomes an extensive thermal-like quantity and is enhanced by heavy-quark backreaction through the monotonic $\rho$-dependence of the coefficients $\widetilde{S}$.
[…]
6 HEN for bipartite subsystems
Having analyzed the adjacent subsystems configuration, we now extend our study to the bipartite case, where the subsystem $A$ and its complement $A^{c}$ together form the entire boundary theory. Unlike the previous setups, this configuration captures the global entanglement structure of the plasma and provides a more direct comparison between quantum and thermal correlations. In particular, the bipartite negativity is expected to be sensitive to the thermal entropy density at high temperature, while at low temperature it should predominantly reflect the underlying vacuum-like entanglement structure. Furthermore, the presence of backreaction introduces nontrivial corrections that allow us to probe how heavy quark density modifies the global entanglement content of the system.
To compute the bipartite HEN, we divide the $A^{c}$ in two parts denoted by $B_{1}$ and $B_{2}$ on the either side of $A$. We can define the regions as,
$B_{1}\in[-\frac{l}{2},-L],\qquad A\in[-\frac{l_{1}}{2},\frac{l_{1}}{2}],\qquad B_{2}\in[\frac{l}{2},L].$ (6.1)
The HEN for this case is given by [Chaturvedi:2016rft]
$\mathcal{E}=\lim_{B\to A^{c}}\biggl[\frac{3}{16G_{N}}(2\mathcal{A}_{A}+\mathcal{A}_{B_{1}}+\mathcal{A}_{B_{2}}-\mathcal{A}_{A\cup B_{1}}-\mathcal{A}_{A\cup B_{2}})\biggr]$ (6.2)
Where $B=B_{1}\cup B_{2}$, we can choose $B_{1}=B_{2}$ then,
$\mathcal{E}=\lim_{B\to A^{c}}\frac{3}{8G_{N}}(\mathcal{A}_{A_{1}}+\mathcal{A}_{B_{1}}-\mathcal{A}_{A\cup B_{1}})$ (6.3)
where $B\to A^{c}$ is the bipartite limit and can be imposed as $L\to\infty$.
6.1 Low temperature analysis
Starting from the general prescription (6.3) and using the low-temperature expansions for the area functionals, one obtains for the bipartite configuration as
$\begin{split}\mathcal{E}&=\lim_{B\to A^{c}}\frac{3}{8G_{N}}\Bigg\{A_{0}+R^{d-1}S_{0}+R^{d-1}\rho\,S_{0}S_{2}\Big(\tfrac{4\pi T_{f}}{d}\Big)^{d-1}\Big[\Big(\tfrac{L}{l}\Big)^{d-2}+\Big(\tfrac{L}{L-\tfrac{l}{2}}\Big)^{d-2}-\Big(\tfrac{L}{L+\tfrac{l}{2}}\Big)^{d-2}\Big]\\ &\qquad\qquad+R^{d-1}S_{0}S_{1}(1-\rho)\Big(\tfrac{4\pi T_{f}}{d}\Big)^{d}\Big[\Big(\tfrac{L}{l}\Big)^{d-2}l^{d}+\Big(\tfrac{L}{L-\tfrac{l}{2}}\Big)^{d-2}\big(L-\tfrac{l}{2}\big)^{d}-\Big(\tfrac{L}{L+\tfrac{l}{2}}\Big)^{d-2}\big(L+\tfrac{l}{2}\big)^{d}\Big]\\ &\qquad\qquad+\mathcal{O}\big((T_{f}l)^{2d-2}\big)\Bigg\}\,.\end{split}$ (6.4)
Taking the large transverse-volume limit $L\to\infty$ (with $l$ held fixed and $L\gg l$) simplifies the geometric prefactors and yields the dominant large–$L$ form
$\begin{split}\mathcal{E}=\frac{3}{8G_{N}}\Bigg\{A_{0}+R^{d-1}S_{0}+R^{d-1}S_{0}L^{d-2}\Big[\,l^{2-d}\;+\;\rho\,S_{2}\,l^{2-d}\Big(\tfrac{4\pi T_{f}}{d}\Big)^{d-1}&\;+\;(1-\rho)\,S_{1}\,l^{4-d}\Big(\tfrac{4\pi T_{f}}{d}\Big)^{d}\;\Big]\\ &+\;\mathcal{O}\big((T_{f}l)^{2d-2}\big)\Bigg\},\end{split}$ (6.5)
where $A_{0}$ is the UV-divergent area-law piece and $S_{0},S_{1},S_{2}$ are $d$-dependent constants defined previously. One can draw the following conclusions from (6.5),
  • Within the finite part the leading geometric scaling is $l^{2-d}$ (the vacuum-like short-distance scaling for a strip), with two distinct backreaction corrections controlled by $\rho$.
  • There are two $\rho$-dependent corrections with different parametric dependence: first is A $\rho$–linear correction. This term enters with the same geometric scaling $l^{2-d}$ as the leading finite term and therefore represents a multiplicative modification of the vacuum-like contribution. The second one is $(1-\rho)$. This carries an extra factor of $l^{2}$ relative to the first correction and is parametrically higher order in $T_{f}$.
  • For the low-temperature expansion and for sufficiently small strip width $l$, the $\rho$–linear term (scaling as $T_{f}^{\,d-1}\,l^{2-d}$) typically dominates over the $(1-\rho)$ term.
This mixed behavior contrasts with the adjacent-interval low-$T_{f}$ result (where a $(1-\rho)T_{f}^{d}$ correction suppressed negativity); the bipartite configuration therefore exhibits a subtler interplay between competing $\rho$-dependent effects. The origin of this difference is geometric: the bipartite construction samples area contributions from large RT surfaces (via the $L^{d-2}$ factor), producing the additional $\rho S_{2}T_{f}^{d-1}$ multiplicative modification which is absent (or subleading) in the adjacent formulae.
6.2 High Temerature analysis
[…]
$\mathcal{E}\;\approx\;\lim_{B\to A^{c}}\frac{3}{8G_{N}}\left\{\frac{3R^{d-1}L^{d-2}}{2G_{N}}\left(\frac{4\pi T_{f}}{d}\right)^{d-2}\widetilde{S}(\rho,d)\right\},$ (6.6)
where $\widetilde{S}(\rho,d)$ is the dimensionless coefficient defined in eq.˜3.14 that increases monotonically with the backreaction parameter $\rho$. The important observation in this regimes are as follows
  • It exhibits the $T_{f}^{\,d-2}$ scaling shown in eq.˜6.6. This reflects the fact that, the dominant contribution to the relevant RT surfaces comes from regions of the bulk close to the horizon, so the negativity inherits the thermal (extensive) scaling of horizon-proximate geometric data.
[…]
The enhancement of $\mathcal{E}$ with $\rho$ at high temperature can be interpreted as follows: adding heavy-quark backreaction modifies the near-horizon geometry so that the bulk area-elements contributing to the negativity increase. Since high temperature already biases the system towards thermal (classical) correlations, the backreaction amplifies the same geometrical measure that feeds into the negativity, producing an overall larger value.
In the next section, we extend this analysis to the case of two disjoint subsystems, where the competition between thermal and quantum correlations manifests through a richer phase structure and provides further insight into the effect of backreaction on entanglement.
7 HEN for two disjoint subsystems
Let us denote two disjoint subsystems by $A_{1}$, $A_{2}$ and the region between them is by $A_{m}$ as follows
$x^{1}\in[-\frac{l_{1}}{2},\frac{l_{1}}{2}],\hskip 14.22636ptx^{i}\in[-\frac{L}{2},\frac{L}{2}]$ (7.1)
$x^{1}\in[-\frac{l_{2}}{2},\frac{l_{2}}{2}],\hskip 14.22636ptx^{i}\in[-\frac{L}{2},\frac{L}{2}]$ (7.2)
$x^{1}\in[-\frac{l_{m}}{2},\frac{l_{m}}{2}],\hskip 14.22636ptx^{i}\in[-\frac{L}{2},\frac{L}{2}]$ (7.3)
To calculate the HEN for two disjoint subsystems $A_{1}$ and $A_{2}$ we follow the the method outlined in [KumarBasak:2020viv]. Lets denotes the turning points by $z_{t_{i}}=z_{t_{1}},z_{t_{2}},z_{t_{m}}$ of the RT surfaces anchored from the boundary of $A_{1}$, $A_{2}$ and $A_{m}$ respectively. The HEN for this disjoint case is given by
$\mathcal{E}=\frac{3}{16G_{N}}(\mathcal{A}_{A_{1}\cup A_{m}}+\mathcal{A}_{A_{m}\cup A_{2}}-\mathcal{A}_{A_{1}\cup A_{m}\cup A_{2}}-\mathcal{A}_{A_{m}})$ (7.4)
Now we will only focus on two regimes as done in previous section.
7.1 Low temperature limit
Following the construction introduced in eq.˜3.11 and eq.˜7.4. At low effective temperature ($T_{f}l_{j}\ll 1$), one obtains the analytic expression for HEN as follow
$\begin{split}\mathcal{E}=\frac{3R^{d-1}L^{d-2}}{16G_{N}}S_{0}\Biggl[&(l_{1}+l_{m})^{2-d}+(l_{2}+l_{m})^{2-d}-(l_{1}+l_{2}+l_{m})^{2-d}-(l_{m})^{2-d}\\ &\hskip 8.5359pt-\,2S_{0}S_{1}(1-\rho)\,l_{1}l_{2}\left(\frac{4\pi T_{f}}{d}\right)^{\!d}+\mathcal{O}\!\left((T_{f}l_{j})^{2d-2}\right)\Biggr],\end{split}$ (7.5)
where $S_{0}$ and $S_{1}$ are $d$-dependent constants previously defined.
It is instructive to verify that in the limit $l_{m}\to 0$, the above expression correctly reproduces the result for the adjacent interval configuration obtained in eq.˜5.4. Indeed, setting $l_{m}=0$ in eq.˜7.5 yields eq.˜5.4 which is precisely the low-temperature adjacent HEN result.
From eq.˜7.5 one can observe that
  • The leading contribution in (7.5) is proportional to $S_{0}$ and depends solely on the geometric configuration of the disjoint intervals. This reflects the universal short-distance entanglement structure of the boundary theory.
  • The backreaction enters explicitly through the correction term proportional to $(1-\rho)\,T_{f}^{\,d}$. Similar to the adjacent case, this term suppresses the negativity, indicating that the presence of heavy quarks diminishes quantum correlations at low temperature by enhancing thermal effects.
  • The separation width $l_{m}$ plays a nontrivial role: as $l_{m}$ increases, the negativity decreases due to reduced overlap of the entanglement wedges, consistent with the expectation that large separations weaken correlations.
In summary, the low-temperature behavior of HEN for disjoint intervals captures both the universal short-distance scaling and the suppressive effect of backreaction. In the next section, we will analyze the high-temperature regime, where the interplay between thermal contributions and $\rho$-dependent corrections leads to qualitatively different behavior.
7.2 High temperature limit
We now turn to the analysis of HEN for two disjoint subsystems in the high-temperature regime under $l_{m}\ll l$ condition. Following the procedure outlined in [KumarBasak:2020viv], and employing the general expressions eqs.˜3.13 and 7.4, one obtains
$\mathcal{E}=\frac{3L^{d-2}R^{d-1}}{8G_{N}}\left(\frac{4\pi T_{f}}{d}\right)^{\!d}\Bigl[\widetilde{S}_{1m}(\rho,d)+\widetilde{S}_{2m}(\rho,d)-\widetilde{S}_{1m2}(\rho,d)\Bigr]-\frac{3}{4}S_{A_{m}}^{low},$ (7.6)
where the coefficients $\widetilde{S}(\rho,d)$ are defined in eq.˜3.14 and $S_{A_{m}}^{\mathrm{low}}$ is given by eq.˜4.29.
As a nontrivial check of consistency, we can take the limit $l_{m}\to 0$ in (7.6), which yields
$\mathcal{E}=\frac{3L^{d-2}R^{d-1}}{8G_{N}}\left(\frac{4\pi T_{f}}{d}\right)^{\!d}\Bigl[\widetilde{S}_{1}(\rho,d)+\widetilde{S}_{2}(\rho,d)-\widetilde{S}_{12}(\rho,d)\Bigr],$ (7.7)
precisely matching the adjacent interval result given in eq.˜5.5. This provides a strong consistency check for the validity of the prescription. The following are the main observations for this case:
  • The negativity scales as $T_{f}^{\,d}$, in contrast to the $T_{f}^{\,d-2}$ scaling observed for the adjacent case. This enhanced scaling originates from the explicit dependence of the disjoint configuration on the separation width $l_{m}$.
  • The quantities $\widetilde{S}_{1m}(\rho,d)$, $\widetilde{S}_{2m}(\rho,d)$, $\widetilde{S}_{1m2}(\rho,d)$, and $\widetilde{S}_{m}(\rho,d)$ capture the geometric and backreaction-dependent contributions. Each of these functions increases monotonically with $\rho$, implying that the negativity is enhanced as the quark density grows. This behavior is consistent with the previous cases at high temperature but is in sharp contrast with the suppression observed in the low-temperature regime.
  • The separation width $l_{m}$ introduces an additional control parameter: increasing $l_{m}$ suppresses the negativity, reflecting the weakening of correlations between spatially separated subsystems. The balance between this geometric suppression and the $\rho$-induced enhancement determines the overall behavior of the negativity in this regime.
The analysis of (7.5) and (7.6) reveals the behavior of HEN for disjoint subsystems across thermal regimes. At low temperature, the leading $\rho$-dependent correction appears at order $T_{f}^{\,d}$ and acts to suppress the negativity, consistent with the intuition that heavy quark backreaction enhances thermal correlations at the cost of quantum correlations. In contrast, at high temperature the negativity scales as $T_{f}^{\,d}$ with coefficients $\widetilde{S}(\rho,d)$ that grow monotonically with $\rho$, resulting in an enhancement of quantum correlations in the thermal phase. The separation width $l_{m}$ further modulates this behavior, reducing correlations as the subsystems are pulled apart, but never altering the qualitative difference between the two thermal regimes. This comparative picture establishes that the influence of backreaction is highly regime-dependent: it diminishes entanglement at low temperature while enhancing it at high temperature.
8 Critical separation and entanglement–disentanglement transition
In this section we will compute the critical separation between two subregion in disjoint configuration. In [Chakrabortty:2020ptb], the critical separation between two subsystems is studied using holographic mutual information. The critical separation is defined as the value of the intermediate width at which the HEN vanishes. The HEN for disjoint subsystems from eq.˜7.4 can be rewritten as,
$\mathcal{E}=\frac{3}{4}\Big(S(l_{1}+l_{m})+S(l_{2}+l_{m})-S(l_{1}+l_{2}+l_{m})-S(l_{m})\Big).$ (8.1)
At critical width $l_{m}=l_{c}$ we have,
$\mathcal{E}(l_{c})=0.$ (8.2)
We focus on the physically relevant regime $l_{c}\ll l_{1},l_{2}$, where the entropies associated with the larger subsystems probe the high-temperature limit, while the entropy corresponding to the separation region remains in the low-temperature regime. Accordingly,
$\begin{split}&S(l_{1}+l_{c})=S^{\text{high}}(l_{1}+l_{c}),\qquad S(l_{2}+l_{c})=S^{\text{high}}(l_{2}+l_{c}),\\ &S(l_{1}+l_{2}+l_{c})=S^{\text{high}}(l_{1}+l_{2}+l_{c}),\qquad S(l_{c})=S^{\text{low}}(l_{c}).\end{split}$ (8.3)
From the low-temperature expansion of the holographic entanglement entropy eq.˜3.11, only the leading non-analytic term $S_{0}\,l^{2-d}$ survives for the narrow separation region $l=l_{c}$, since all analytic thermal corrections are parametrically suppressed for $l_{c}T_{f}\ll 1$. Moreover, the UV-divergent term cancels identically in the negativity combination. Therefore, to leading order,
$S^{\text{low}}(l_{c})=\frac{R^{d-1}L^{d-2}}{2G_{N}}\Big[S_{0}\,l_{c}^{2-d}+\mathcal{O}(l_{c}T_{f}^{d-1})\Big].$ (8.4)
The high-temperature HEE for a strip of width $l$ is given by eq.˜3.13 and can be rewritten as,
$S^{\text{high}}(l)=\frac{R^{d-1}L^{d-2}}{2G_{N}}\Big[\alpha_{d}\,T_{f}^{d-2}\,l+\beta_{d}(\rho)\,T_{f}^{d-2}\Big],$ (8.5)
with the effective coefficients
$\alpha_{d}=\frac{1}{2}\left(\frac{4\pi}{d}\right)^{d-1},$ (8.6)
$\beta_{d}(\rho)=\widetilde{S}(\rho,d)\left(\frac{4\pi}{d}\right)^{d-2}.$ (8.7)
The coefficient $\alpha_{d}$ controls the universal extensive thermal contribution, while $\beta_{d}(\rho)$ captures the finite backreaction-dependent correction arising from the near-horizon geometry. Using (8.5), the combination of high-temperature entropies appearing in (8.1) simplifies to
$\begin{split}&S^{\text{high}}(l_{1}+l_{c})+S^{\text{high}}(l_{2}+l_{c})-S^{\text{high}}(l_{1}+l_{2}+l_{c})\\ &\qquad=\frac{R^{d-1}L^{d-2}}{2G_{N}}\Big[\alpha_{d}\,T_{f}^{d-2}\,l_{c}+\beta_{d}(\rho)\,T_{f}^{d-2}\Big],\end{split}$ (8.8)
where all dependence on $l_{1}$ and $l_{2}$ cancels exactly. Imposing the condition (8.2) and combining (8.4) with (8.8), we obtain the algebraic equation
$\alpha_{d}\,T_{f}^{d-2}\,l_{c}+\beta_{d}(\rho)\,T_{f}^{d-2}=S_{0}\,l_{c}^{2-d}\left[1+\mathcal{O}(l_{c}T_{f})\right].$ (8.9)
In the regime $l_{c}T_{f}\ll 1$, the power-law term dominates, yielding the leading-order solution
$l^{(0)}_{c}=\left(\frac{S_{0}}{\beta_{d}(\rho)\,T_{f}^{d-2}}\right)^{\!\frac{1}{d-2}}.$ (8.10)
Since $\beta_{d}(\rho)$ increases monotonically with the backreaction parameter $\rho$, the critical separation decreases as $\rho$ increases. This indicates that backreaction suppresses long-range distillable entanglement, leading to an earlier onset of the entanglement–disentanglement transition. This behavior is in sharp contrast with the enhancement observed for the EWCS and reflected entropy at low temperature, underscoring the distinct physical information captured by holographic entanglement negativity.
In [Chakrabortty:2020ptb], the critical separation between two disjoint strips was analyzed using holographic mutual information (HMI), where the transition point corresponds to a change in the connectivity of the entanglement wedge. The resulting critical distance reflects the persistence of total correlations, including both quantum and classical contributions. In contrast, the critical separation $l_{c}$ obtained from HEN characterizes the vanishing of distillable quantum entanglement between the subsystems. As a consequence, the negativity-based critical distance is generically smaller than that obtained from HMI. This hierarchy is physically expected: while classical correlations can survive at larger separations and temperatures, quantum entanglement is more fragile and is suppressed more efficiently by thermal and backreaction effects. In the present backreacted AdS black hole background, increasing the quark density parameter $\rho$ enhances the thermal entropy and accelerates the disentanglement of quantum correlations, leading to a monotonic decrease of $l_{c}$ from negativity. On the other hand, the HMI-based critical distance exhibits a comparatively weaker suppression with $\rho$, consistent with its sensitivity to classical correlations encoded in the geometry of the entanglement wedge. The comparison thus highlights the complementary roles of HMI and HEN: while both diagnose correlation phase transitions, negativity provides a stricter and more refined probe of quantum entanglement in strongly coupled plasmas.
9 Summary and Discussion
[…]
Overall, the observations made here provide a deeper understanding of entanglement and related phenomena in strongly coupled conformal field theories with backreaction at finite temperature. These results establish that the influence of heavy flavour backreaction is not universal but strongly regime dependent as it reduces entanglement in the vacuum like phase and enhances it in the thermal phase. This dual role highlights HEN as a sharper diagnostic of the balance between quantum and thermal correlations than HEE alone. In our study of entangled-disentangled phase transition using HEN for disjoint configuration we found that the critical separation decreases with backreaction. Also the critical separation from HEN ($l_{c}^{HEN}$) is always smaller then obtained from HMI ($l_{c}^{HMI}$) i.e. $l_{c}^{HEN}<l_{c}^{HMI}$.
Future directions include performing a systematic numerical study in the intermediate regime $T_{f}l\sim\mathcal{O}(1)$ to map out crossover behaviour, exploring time-dependent quenches to connect with thermalization dynamics and extending the analysis to higher-derivative gravity or string-theoretic embeddings to quantify $1/N$ and $\alpha^{\prime}$ corrections. Such studies could shed further light on the microscopic structure of entanglement in strongly coupled plasmas with flavour backreaction and on the universal features of entanglement negativity in holographic duals.
Acknowledgments
S.P would like to thank Prof. Hemwati Nandan for his support during the visit to HNB Garhwal University. H.P acknowledges the support of this work by NCTS.
Appendix A Numerical coefficients
The constants used in eq.˜4.46 are given by
$A_{0}$ $=\sqrt{\pi}\,\frac{\Gamma\!\left(-\tfrac{1}{4}\right)}{2\,\Gamma\!\left(\tfrac{1}{4}\right)}+W_{a,0}+\frac{\pi^{2}}{9\sqrt{3}},\quad A_{1}=A_{0}-\,W_{a,1}-\frac{2}{\sqrt{3}}+\frac{2\ln 3}{\sqrt{3}}$ (A.1)
$D_{0}$ $=3\,W_{b,0}-3\,W_{c,0}+3\big(G_{1}-G_{2}\big),\quad G_{1}=\frac{\Gamma(\tfrac{3}{2})\Gamma(\tfrac{5}{4})}{\Gamma(\tfrac{7}{4})},\quad G_{2}=\frac{(\Gamma(\tfrac{3}{2}))^{2}}{2}$
$D_{1}$ $=D_{0}-3\,W_{b,1}+3\,W_{c,1}+3(-2G_{1}+3G_{2})-\frac{2\ln 3}{\sqrt{3}}$
$W_{a,0}$ $=\sum_{n=1}^{\infty}\tilde{a}_{n},\quad W_{a,1}=\sum_{n=1}^{\infty}(3n)\,\tilde{a}_{n},\quad W_{b,0}=\sum_{n=1}^{\infty}\tilde{b}_{n},\quad W_{b,1}=\sum_{n=1}^{\infty}(3n+2)\,\tilde{b}_{n},$
$W_{c,0}$ $=\sum_{n=1}^{\infty}\tilde{c}_{n},\quad W_{c,1}=\sum_{n=1}^{\infty}(3n+3)\,\tilde{c}_{n}$
$\tilde{a}_{n}$ $=\frac{1}{3n-1}\frac{\Gamma(n+\tfrac{1}{2})}{\Gamma(n+1)}\frac{\Gamma\!\left(\tfrac{3n+3}{4}\right)}{\Gamma\!\left(\tfrac{3n+5}{4}\right)}-\frac{2}{3\sqrt{3}\,n^{2}},\quad\tilde{b}_{n}=\frac{1}{3n+1}\frac{\Gamma(n+\tfrac{3}{2})}{\Gamma(n+1)}\frac{\Gamma\!\left(\tfrac{3n+5}{4}\right)}{\Gamma\!\left(\tfrac{3n+7}{4}\right)}-\frac{2}{3\sqrt{3}\,n},$
$\tilde{c}_{n}$ $=\frac{1}{3n+2}\frac{\Gamma(n+\tfrac{3}{2})}{\Gamma(n+1)}\frac{\Gamma\!\left(\tfrac{3n+6}{4}\right)}{\Gamma\!\left(\tfrac{3n+8}{4}\right)}-\frac{2}{3\sqrt{3}\,n}.$